### v2

# MNSC-0001-2024.00

[…]
MNSC-0001-2024.00
[…]
Banerjee, Hssaine, and Sinclair
[…]
Online Fair Allocation of Perishable Resources
[…]
Online Fair Allocation of Perishable Resources
[…]
Sid Banerjee
[…]
Chamsi Hssaine
[…]
Sean R. Sinclair
[…]
We consider a practically motivated variant of the canonical online fair allocation problem: a decision-maker has a budget of perishable resources to allocate over a fixed number of rounds. Each round sees a random number of arrivals, and the decision-maker must commit to an allocation for these individuals before moving on to the next round. The goal is to construct a sequence of allocations that is envy-free and efficient. Our work makes two important contributions toward this problem: we first derive strong lower bounds on the optimal envy-efficiency trade-off, demonstrating that a decision-maker is fundamentally limited in what she can hope to achieve relative to the no-perishing setting; we then design an algorithm achieving these lower bounds which takes as input $(i)$ a prediction of the perishing order, and $(ii)$ a desired bound on envy. Given the remaining budget in each period, the algorithm uses forecasts of future demand and perishing to adaptively choose from one of two carefully constructed guardrail quantities. We demonstrate our algorithm’s strong numerical performance — and state-of-the-art, perishing-agnostic algorithms’ inefficacy — on simulations calibrated to a real-world dataset.
[…]
Online Resource Allocation, Fairness, Perishability
1 Introduction
Despite a consistent decline in food insecurity in the United States over the past decade, 2022 saw a marked increase in individuals struggling to access enough food to fulfill basic needs. A recent report by the U.S. Department of Agriculture found that over 44 million individuals faced some form of hunger in 2022 — 45% more than the previous year (Godoy 2023). Due in part to rising food prices and the rolling back of pandemic-era social security measures, this disturbing statistic has further underscored the important role of local food banks; for instance, the Feeding America network of food banks, food agencies, and local food programs distributed over 5.2 billion meals that same year (Feeding America 2022).
In distributing food throughout their operating horizon, food banks have two competing objectives: distributing as much food as possible to communities in need, and ensuring equitable access to donations. This tension has attracted much attention in the operations literature, with recent work characterizing the fundamental trade-offs between fairness and efficiency in sequential allocation problems (Bertsimas et al. 2011, Donahue and Kleinberg 2020, Lien et al. 2014, Manshadi et al. 2023, Sinclair et al. 2023, Onyeze et al. 2025b). Understanding such trade-offs in theory is useful, as they allow a system designer to recognize and choose their desired operating point, balancing the loss in efficiency and equity. Despite the useful insights derived in prior work, to the best of our knowledge an important reality of food bank operations remains overlooked: waste due to spoilage of perishable goods.
Perishable goods constitute a substantial portion of food donations. For instance, 35% of the food distributed annually by the Food Bank of Central and Eastern North Carolina is perishable (Sengul Orgut and Lodree 2023). Similarly, in 2024 alone, the Los Angeles Regional Food Bank distributed over 38 million pounds of fresh produce, representing 29% of the total amount of food distributed throughout the year (Los Angeles Regional Food Bank 2025). At the same time, the effective distribution of perishable goods has long been a significant challenge for these organizations. For example, in 2020 Feeding America reported that 3% of the food received by its partner agencies was disposed of due to spoilage. Similar percentages were reported by other food rescue organizations in the United States, with Boston- and Rochester-based organizations respectively reporting that 2.4% and 6% of all received food was wasted due to spoilage (U.S. Environmental Protection Agency 2020).
These figures highlight a persistent operational gap between the supply of perishable goods and their successful distribution. The reason for this gap is two-fold. First, food banks lack control over the kinds of donated items, meaning that the supply of goods that are about to perish can far exceed the demand for these goods. Ohls et al. (2002) highlighted this as a significant concern: in interviews with over 300 food banks, they found that around 20% of food banks distribute less than 90% of the food they receive, partly due to this issue. Second, even when this is not the case, true expiration dates are frequently unknown. Rather, product packaging commonly indicates “best-by,” “use-by,” or “sell-by” dates, after which products are still safe to consume (U.S. Department of Agriculture, Food Safety and Inspection Service 2019). So, while these dates may provide an informative signal as to when goods will perish (e.g., one week after the “best-by” date), the actual time at which this occurs is not known ex-ante. This problem is further exacerbated within the context of food banks, where individuals are encouraged to donate perishables that are past their “best-by” dates (City Harvest 2019, WQAD News 8 2025). Not only does this reduce the amount of time food banks have to allocate these items before they spoil, but it also creates significant uncertainty in the quality of the donations (Zou et al. 2025).
Intuitively, the equity-efficiency trade-off should be exacerbated in the presence of perishables: while equity requires a decision-maker to allocate conservatively to preserve resources for future demand (Sinclair et al. 2023), perishability starts a “race against time.” As goods perish due to a slow allocation rate, not only is efficiency harmed, but so may be equity, since spoilage runs the risk of a decision-maker running out of items, with nothing left to give out by the end of the operating horizon. Thus motivated, this paper seeks to answer the following questions:
Do established equity-efficiency trade-offs in dynamic environments persist in the presence of perishable goods? If not, what limits do they impose on fair and efficient allocation? Can we design policies that perform well under these limits?
1.1 Our Contributions
We consider a model in which a decision-maker has a fixed budget of items (also referred to as goods, or resources) to be allocated over $T$ discrete rounds. An a priori unknown number of individuals arrives in each period, each seeking a share of goods. Each agent is characterized by an observable type (drawn from a known, potentially time-varying distribution), associated with a linear utility function over the received allocation. Moreover, each unit of good has a stochastic perishing time (similarly drawn from a known distribution, independent of incoming arrivals), after which the good spoils and can no longer be allocated. The decision-maker allocates goods according to a fixed ordering (also referred to as perishing prediction, or allocation schedule). For instance, if each good has an associated “best-by” date, in practice the decision-maker may allocate the goods in increasing order of these dates. The goal is to find a policy that trades off between three ex-post metrics:
  1. 1.
    Hindsight Envy (Envy) – Maximum difference in utility obtained by any two agents.
  2. 2.
    Counterfactual Envy $(\Delta_{\text{\it EF}})$ – Maximum difference between the utility obtained by any agent, and their utility under the static, proportional allocation (optimal in the no-perishing setting).
  3. 3.
    Inefficiency $(\Delta_{\text{\it efficiency}})$ – Amount of unallocated (including spoiled) goods at the end of the horizon.
For this setting, we first characterize the fundamental limits that perishability places on any online algorithm. To see the considerable impact that perishability can have on counterfactual envy, consider an extreme scenario in which all items perish at the end of the first round. Clearly, there is no hope of achieving low envy in such a setting, since future demand can never be satisfied. Our first main contribution identifies a necessary and sufficient condition on the joint perishing and arrival distribution for low counterfactual envy to be an achievable desideratum (Theorem 3.3). From a managerial perspective, this characterization — which at a high level states that the cumulative perishing must lag behind cumulative arrivals — can be viewed as guidance on the composition of perishable donations that the decision-maker should accept. It moreover underscores the importance of leveraging joint information over perishing and demand: if demand is back-loaded, perishing times should be concentrated late in the horizon; if most demand arrives early, however, early perishing times are acceptable.
For this class of processes, which we term offset-expiring, zero spoilage occurs if goods are allocated in the exact order in which they perish. Due to stochasticity in goods’ perishing times, however, the decision-maker’s allocation schedule may be misaligned with the realized perishing times of goods. For instance, while one good may have an earlier “best-by” date than another, the latter good may spoil earlier. Alternatively, if the decision-maker allocates goods in increasing order of expected perishing time, the variance of the perishing times may cause goods with later expected perishing times to perish earlier than others. We show that the inherent aggressiveness of the perishing process (i.e., in cases where the process is not offset-expiring) and inaccuracies in the allocation schedule pose an insurmountable barrier to any algorithm’s performance by deriving lower bounds on any algorithm’s envy and inefficiency. These lower bounds are parameterized by an unavoidable loss due to spoilage, also referred to as the perishing-induced loss and denoted by $\mathcal{L}^{\textsf{perish}}$ (Definition 3.7), implicitly characterized as the solution to a fixed-point equation (Theorem 3.8). In contrast to the no-perishing setting, in which the only source of loss is exogenous uncertainty in demand, in our setting the loss due to spoilage is endogenous: it crucially depends on the rate at which the algorithm allocates items. This endogeneity poses significant challenges in the analysis of any adaptive algorithm; designing a tractable approach to analyzing ex-post spoilage is the main technical contribution of our work. Additionally, the lower bounds we derive give rise to the key insight that, contrary to the “race against time” intuition under which a decision-maker must increase the allocation rate to prevent avoidable spoilage, achieving low hindsight envy in the presence of unavoidable spoilage requires a decision-maker to potentially allocate significantly less than the proportional allocation. Hence, perishability throws a wrench into the well-studied envy-efficiency trade-off: while hindsight and counterfactual envy are aligned in the no-perishing setting, these two may be at odds when goods spoil, since only high counterfactual-envy solutions may yield low hindsight-envy. The tension between efficiency and equity is exacerbated for the same reason, relative to the classical setting.
In our final technical contribution, we leverage these insights to construct an adaptive threshold algorithm (Algorithm 1) that achieves these lower bounds (Theorem 4.1). Our algorithm takes as input $(i)$ the fixed allocation schedule, $(ii)$ a desired upper bound on hindsight envy $L_{T}$, and $(iii)$ a high-probability parameter $\delta$. Given these inputs, it computes a high-probability lower bound on a budget-respecting zero-hindsight-envy solution, and an “aggressive” efficiency-improving allocation that is $L_{T}$ away. In each round the algorithm chooses which of the two quantities to allocate to each individual, cautiously doing so by constructing pessimistic forecasts of future arrivals and perishing. While this algorithm is similar in flavor to state-of-the-art algorithms for the no-perishing setting (Sinclair et al. 2023), the main challenge it contends with is forecasting endogenous future spoilage. We overcome this challenge by leveraging an exogenous “slow” consumption process that tractably decouples past allocations from future perishing. Crucially, we show that future spoilage under this slow process upper bounds our algorithm’s future spoilage.
[…]
Figure 1: Graphical representation of Theorems 4.1 and 3.8, illustrating the envy-efficiency trade-off ($\Delta_{\text{\it efficiency}}$ vs. $\Delta_{\text{\it EF}}$) achieved by Perishing-Guardrail (Algorithm 1) in the “high-envy, low-perishing” regime. Here, $\mathcal{L}^{\textsf{perish}}$ is used to denoted the unavoidable perishing-induced loss (Definition 3.7). The vertical and horizontal dotted lines represent the impossibility results due to demand and perishing uncertainty, respectively. The region below the solid line represents the impossibility due to the envy-efficiency trade-off; the green region is the achievable region for Perishing-Guardrail.
Our bounds give rise to three salient regimes that dictate the extent to which a decision-maker is restricted in leveraging inequity to improve efficiency (and vice versa). These three regimes are a function of both $L_{T}$ and the unavoidable perishing-induced loss dictating our lower bounds, and at a high level are as follows:
  1. 1.
    Low-envy: There are no efficiency gains from deviating from the equitable solution, where all individuals receive the same amount.
  2. 2.
    High-envy, high-perishing: Similarly here, inefficiency is invariant to $L_{T}$.
  3. 3.
    High-envy, low-perishing: Inefficiency decreases as $1/L_{T}$, until reaching the unavoidable perishing-induced loss throughout the horizon, $T\mathcal{L}^{\textsf{perish}}$. This regime is represented pictorially in Figure 1.
We conclude our theoretical results by providing explicit characterizations of the unavoidable perishing-induced loss for important special cases (Propositions 4.16, 4.17, 4.18 and 4.19). These characterizations shed light on the salient drivers of this loss: $(i)$ how aggressively goods perish relative to the rate at which they are demanded, and $(ii)$ whether the allocation schedule approximately preserves the order induced by the true perishing times. Our results reveal that the variability of the perishing distribution plays a crucial role in both cases.
We complement our theoretical bounds by testing the practical performance of our algorithm on both synthetic and real-world datasets. Our experiments show that the unfairness required to achieve efficiency gains is order-wise larger than in settings without perishable resources. Additionally, they underscore the weakness of perishing-agnostic online algorithms. We observe that these latter algorithms are incapable of leveraging unfairness to improve efficiency across a variety of perishing regimes. In contrast to these, our algorithm’s construction of a perishing-aware baseline allocation is necessary to mitigate stockouts across all — rather than simply worst-case — instances. These include instances where offset-expiry fails to hold with high probability, as is the case in the real-world dataset we use to calibrate our experiments (Keskin et al. 2022). Perhaps most surprisingly, despite our baseline allocation being significantly lower than that of algorithms that don’t take into account perishability, our algorithm is more efficient than these more aggressive algorithms, in addition to being more fair. This observation contradicts the “race against time” intuition that aggressive allocations are necessarily more efficient than cautious ones. Finally, we numerically explore the question of selecting practical allocation schedules, with the goal of minimizing the unavoidable loss due to perishing. Our main managerial insight is that ordering items according to a high-probability lower bound on their perishing time is a robust, interpretable choice for decision-makers; it naturally takes into account items’ expected perishing times, all the while hedging against the variability of the perishing process.
Finally, we note that while our work is motivated by food bank operations, perishability is also an important consideration in medical applications, such as blood bank operations (Bar-Lev et al. 2017) and vaccine distribution (Manshadi et al. 2023). For these applications, our modeling framework is applicable to settings in which allocation occurs in rural or remote areas, where uncertainty around storage conditions (e.g., temperature and humidity) may affect the lifetime of these goods.
Paper organization.
We next survey the related literature. We present the model in Section 2, and formalize the limits of perishability in Section 3. In Section 4 we design and analyze an algorithm that achieves the derived lower bounds, illustrating the impact of the quality of the allocation schedule and the aggressiveness of the perishing process for important special cases in Section 4.4. We conclude with numerical experiments in Section 5.
1.2 Related Work
Fairness in resource allocation has a long history in the economics and computation literature, beginning with Varian’s seminal work (Varian 1974, 1976). More recently, there has been work studying fairness in operations, including resource-level fairness through load balancing (Chen et al. 2022, Hssaine et al. 2024, Balseiro et al. 2025, Freund et al. 2025), pricing (Cohen et al. 2022, den Boer et al. 2022), incentive design (Freund and Hssaine 2025, Hssaine et al. 2025), algorithmic hiring and online selection (Salem and Gupta 2024, Hu et al. 2025), and societal systems more generally (Gupta and Kamble 2021, Lyu et al. 2025). We highlight the most closely related works below, especially as they relate to online fair allocation; see Aleksandrov and Walsh (2019b) for a survey.
Fair allocation without perishable resources.
There exists a long line of work in which non-perishable resources become available to the decision-maker online, whereas agents are fixed (Benade et al. 2018, Aleksandrov et al. 2015, Mattei et al. 2017, 2018, Aleksandrov and Walsh 2019a, Banerjee et al. 2022, Bansal et al. 2020, Bogomolnaia et al. 2022, He et al. 2019, Aziz et al. 2016, Zeng and Psomas 2020). These models lie in contrast to the one we consider, wherein resources are fixed and individuals arrive online. Papers that consider this latter setting include Kalinowski et al. (2013), which studies welfare maximization with indivisible goods; Gerding et al. (2019) considers a scheduling setting wherein agents have fixed and known arrival and departure times, as well as demand for the resource; Hassanzadeh et al. (2023) allows individuals to arrive in multiple timesteps. A series of papers also consider the problem of fair allocation with minimal disruptions relative to previous allocations, as measured by a fairness ratio, a competitive ratio analog of counterfactual envy in our setting (Friedman et al. 2017, Cole et al. 2013, Friedman et al. 2015). Other works design algorithms with attractive competitive ratios with respect to Nash Social Welfare (Azar et al. 2010, Banerjee et al. 2022) and the max-min objective (Lien et al. 2014, Manshadi et al. 2023). Ma et al. (2025) recently showed how forward-backwards contention resolution schemes can be leveraged to obtain competitive ratios for online rationing in non-perishable settings.
The above papers situate themselves within the adversarial, or worst-case, tradition. A separate line of work considers fair resource allocation in stochastic settings (Donahue and Kleinberg 2020, Elzayn et al. 2019), as we do. The algorithms developed in these papers, however, are non-adaptive: they decide on the entire allocation upfront, before observing any of the realized demand. In contrast, we consider a model where the decision-maker makes the allocation decision in each round after observing the number of arrivals. Freeman et al. (2017) considers a problem in which agents’ utilities are realized from an unknown distribution, and the budget resets in each round. They present algorithms for Nash social welfare maximization and discuss some of their properties. Our work is most closely related to (and indeed, builds upon) Sinclair et al. (2023), which first introduced the envy-efficiency tradeoff we are interested in. Our work considers the more challenging setting of perishable goods, which none of the aforementioned works consider.
Finally, we note the existence of a rich line of work that takes a mechanism design approach to fair division, dating back to the 1940s (Steinhaus 1948). We refer the reader to Moulin (2004), Chapters 11-13 of Brandt et al. (2016), and Chapter 11 of Kroer (2026) for textbook treatments of the topic. Generally speaking, the main objective of these works is the design of nonmonetary mechanisms that fairly allocate a set of resources amongst strategic agents with private preferences for these goods. Much of the work has focused on one-shot (i.e., static) settings, for which there exist strong impossibility results precluding the existence of fair, efficient, and incentive compatible mechanisms (Budish 2011). Notable works that derive approximately fair and incentive-compatible mechanisms for these settings include Budish (2011) and Brânzei et al. (2017). In dynamic settings, models of fair allocation under incentives involve the same set of agents competing for resources across time, which differs from the arrival model we consider here. Recent works that consider this setting include Onyeze et al. (2025a), who study equilibrium existence, and Fikioris et al. (2024), who propose a generalization of the popular dynamic max-min fair allocation mechanism to achieve strong fairness and approximate incentive compatibility guarantees.
We conclude by highlighting the closely related work of Prendergast (2022), which empirically studies the virtual currency-based auction mechanism used by Feeding America to allocate food to its members. This work finds that, compared to a previous queuing mechanism under which food banks were being allocated an equal amount of random food through time, the auction mechanism allows food banks to express preferences across different types of food. As it relates to our work, it would be interesting and practically relevant to study how $(i)$ the timing of allocations, and $(ii)$ varying preferences across resources induce strategic behavior. We leave this for future work.
Dynamic allocation of perishable resources.
Though dynamic resource allocation of perishable goods has a long history in the operations research literature (see, e.g., Nahmias (2011) for a comprehensive survey of earlier literature), to the best of our knowledge, the question of fairly allocating perishable goods has attracted relatively little attention. We highlight the few relevant papers below. Perry (1999) and Hanukov et al. (2020) analyze FIFO-style policies for efficiency maximization in inventory models with Poisson demand and deterministic or Poisson perishing times. Motivated by the problem of electric vehicle charging, Gerding et al. (2019) consider an online scheduling problem where agents arrive and compete for a perishable resource that spoils at the end of every period, and as a result must be allocated at every time step. They consider a range of objectives, including: maximum amount of resources allocated, maximum number of satisfied agents, as well as envy-freeness. Bateni et al. (2022) similarly consider a setting wherein an arriving stream of goods perish immediately. Recent empirical work by Sengul Orgut and Lodree (2023) considers the problem of a food bank equitably and efficiently allocating perishable goods under complete information. Their case study on data from a partnering food bank numerically validates our theoretical results: in low-budget settings, there is little or no benefit to increasing inequity to counteract the inefficiency due (in part) to spoilage. In contrast to these latter papers, the model we consider locates itself within the smaller category of inventory models in which products have random lifetimes. The majority of these assume that items have exponentially distributed or geometric lifetimes (Bakker et al. 2012).
2 Preliminaries
We consider a decision-maker who, over $T$ rounds, must allocate $B$ divisible units (also referred to as items) of a single type of resource among a population of individuals. Let $\mathcal{B}$ denote the set of these $B$ units.
2.1 Basic Setup
Demand model.
At the start of each round $t\in[T]$, a random number of individuals arrives, each requesting a share of units. Each individual is characterized by her type $\theta\in\Theta$, with $|\Theta|<\infty$. Each type $\theta$ is associated with a known utility function $u_{\theta}(x)=w_{\theta}\cdot x$ for a given allocation $x\in\mathbb{R}_{+}$ of the resource, with $w_{\theta}>0$. We let $N_{t,\theta}$ denote the number of type $\theta$ arrivals in round $t$; $N_{t,\theta}$ is drawn independently from a known distribution, with $N_{t,\theta}\geq 1$ almost surely for all $t\in[T],\theta\in\Theta$. This latter assumption is for ease of exposition; our results continue to hold (up to constants) as long as $\mathbb{P}(N_{t,\theta}=0)$ does not scale with $T$. For ease of notation we define $N_{t}=\sum_{\theta\in\Theta}N_{t,\theta}$ and $N=\sum_{t\in[T],\theta\in\Theta}N_{t,\theta}$. We assume $\mathbb{E}[N]=\Theta(T)$, and define $\beta_{avg}=B/\mathbb{E}[N]$ to be the average number of units per individual, with $\beta_{avg}=\Theta(1)$.
Perishing model.
Each unit of resource $b\in\mathcal{B}$ is associated with a perishing time $T_{b}\in\mathbb{N}^{+}$ drawn from a known distribution. Items’ perishing times are independent of one another and of the arrival process, and perishing occurs at the end of each round, after items have been allocated to individuals. For $t\in[T]$, we let $P_{t}=\sum_{b\in\mathcal{B}}\mathds{1}\left\{T_{b}=t\right\}$ denote the number of units of resource perishing in period $t$.
The decision-maker has access to a predicted ordering according to which items perish; we will often refer to this ordering as the allocation schedule. For example, the decision-maker may order items in increasing order of their expected perishing times, or according to a lower confidence bound on their perishing times. As alluded to in the introduction, within the context of food bank operations, the decision-maker may have access to “best-by” dates specified on food packaging and choose to allocate items with earlier “best-by” dates first. We use $\sigma:\mathcal{B}\rightarrow[B]$ to denote this ordering, i.e., $\sigma(b)$ is rank of $b$ in this ordering. For $b\in[B]$, $\sigma^{-1}(b)$ is used to denote the identity of the $b$th ranked item in $\sigma$, with $\sigma^{-1}(1)$ being the item that comes first in the allocation schedule.
Example 2.1
Consider an instance for which $\mathcal{B}=\{\texttt{a},\texttt{b},\texttt{c}\}$, and let $\sigma$ be such that $\sigma(\texttt{a})=2$, $\sigma(\texttt{b})=3$, and $\sigma(\texttt{c})=1$. Then, $\sigma^{-1}(1)=\texttt{c}$, $\sigma^{-1}(2)=\texttt{a}$, and $\sigma^{-1}(3)=\texttt{b}$.
Remark 2.2
In this paper we restrict our attention to static (rather than time-varying and sample path-dependent) allocation schedules, given their practical relevance to the motivating real-world applications described in Section 1. We leave the nonstationary extension to future work.
Additional notation.
For any time-dependent quantity $Y_{t}$, we define $Y_{\leq t}=\sum_{t^{\prime}\leq t}Y_{t^{\prime}}$, $Y_{\geq t}=\sum_{t^{\prime}\geq t}Y_{t^{\prime}}$, along with their strict analogs. We let $w_{max}=\max_{\theta}w_{\theta}$, $\sigma_{t,\theta}^{2}=\text{Var}(N_{t,\theta})<\infty$, and assume $\rho_{t,\theta}=|N_{t,\theta}-\mathbb{E}[N_{t,\theta}]|<\infty$ almost surely. Finally, let $\mu_{\max}=\max_{t}\mathbb{E}\left[N_{t}\right],\sigma^{2}_{\min}=\min_{t,\theta}\sigma^{2}_{t,\theta},\sigma^{2}_{\max}=\max_{t,\theta}\sigma^{2}_{t,\theta}$, and $\rho_{\max}=\max_{t,\theta}\rho_{t,\theta}$. We use $\lesssim$ and $\gtrsim$ to denote the fact that inequalities hold up to polynomial factors of $\beta_{avg},|\Theta|,w_{max},\mu_{\max},\sigma^{2}_{\min},\sigma^{2}_{\max}$, $\log T$, and $\log(1/\delta)$.
2.2 Notions of Fairness and Efficiency
The goal is to design a fair and efficient online algorithm that determines the amount to allocate to all $N_{t,\theta}$ individuals in each round $t$, for all $\theta\in\Theta$, given the remaining budget in each round. We assume this amount is allocated uniformly across all $N_{t,\theta}$ individuals of type $\theta$. We use $X_{t,\theta}^{alg}\in\mathbb{R}_{+}$ to denote the per-individual amount distributed in period $t$, with $X^{alg}=(X_{t,\theta}^{alg})_{t\in[T]}$. Finally, we let $B_{t}^{alg}$ denote the remaining budget at the beginning of period $t$, defined inductively as follows:
[…]
where $\textsc{PUA}_{t}^{alg}$ denotes the quantity of unallocated items that perish at the end of round $t$ (i.e., the perished and unallocated items). We moreover let $\mathcal{B}_{t}^{alg}$ be the set of items remaining at the beginning of period $t$.
Our notions of online fairness and efficiency are motivated by the offline notion of Varian Fairness (Varian 1974), and are the same as those considered in past works (Sinclair et al. 2023).
Definition 2.3 (Counterfactual Envy, Hindsight Envy, and Efficiency)
Given budget $B$, realized demands $(N_{t,\theta})_{t\in[T],\theta\in\Theta}$, perishing times $(T_{b})_{b\in[B]}$, and allocation schedule $\sigma$, for any online allocation defined by $X^{alg}$ we define:
$\bullet\ \text{\emph{Counterfactual Envy}:}$ $\Delta_{\text{\it EF}}$ $\triangleq\max_{t\in[T],\theta\in\Theta}\left|w_{\theta}\left(X_{t,\theta}^{alg}-\frac{B}{N}\right)\right|$ (2)
$\bullet\ \text{\emph{Hindsight Envy}:}$ Envy $\triangleq\max_{t,t^{\prime}\in[T]^{2},\theta,\theta^{\prime}\in\Theta^{2}}w_{\theta}(X_{t^{\prime},\theta^{\prime}}^{alg}-X_{t,\theta}^{alg})$ (3)
$\bullet\ \text{\emph{Inefficiency}:}$ $\Delta_{\text{\it efficiency}}$ $\triangleq B-\sum_{t\in[T],\theta\in\Theta}N_{t,\theta}X_{t,\theta}^{alg}$ (4)
In the offline setting without perishable goods, Varian (1974) established that $X^{opt}_{t,\theta}=B/N$ (referred to as the proportional allocation) is the optimal fair and efficient per-individual allocation. Hence, counterfactual envy $\Delta_{EF}$ can be interpreted as a form of regret with respect to this strong no-perishing benchmark, and can be used to characterize the impact of perishability on our algorithm’s performance. Hindsight envy, on the other hand, measures how differently the online algorithm treats any two individuals across time. Finally, the efficiency of the online algorithm, $\Delta_{\text{\it efficiency}}$, measures how wasteful the algorithm was in hindsight. This could happen in two ways: either through spoilage, or because the decision-maker allocated too conservatively throughout the horizon, thus leaving a large number of unspoiled goods unallocated by $T$.
Even in simple settings without perishability, it is known that these metrics are at odds with each other in online settings. To see this, consider the following two scenarios. On the one hand, an algorithm can trivially achieve a hindsight envy of zero by allocating nothing to individuals in any round; this, however, would result in both high counterfactual envy, in addition to maximal inefficiency. On the other hand, a distance to efficiency of zero can trivially be satisfied by exhausting the budget in the first round, at a cost of maximal hindsight envy as individuals arriving at later rounds receive nothing. Sinclair et al. (2023) formalized this tension for the additive utility setting without perishable resources via the following lower bounds.
Theorem 2.4 (Theorems 1 and 2, Sinclair et al. (2023))
Under any arrival distribution satisfying mild regularity conditions, there exists a problem instance without perishing, such that any algorithm must incur $\Delta_{\text{\it EF}}\gtrsim\frac{1}{\sqrt{T}}$. Moreover, any algorithm that achieves $\Delta_{\text{\it EF}}=L_{T}=o(1)$ or $\textsc{Envy}=L_{T}=o(1)$ must also incur $\Delta_{\text{\it efficiency}}\gtrsim\min\{\sqrt{T},1/L_{T}\}.$
Since settings without perishable resources are a special case of our setting (e.g., a perishing process with $T_{b}>T$ a.s., for all $b\in\mathcal{B}$), this lower bound holds in our case; the goal then is to design algorithms that achieve this lower bound with high-probability. However, as we will see in Section 3, perishing stochasticity is fundamentally distinct from, and more challenging than, demand stochasticity. This difference is particularly salient in regards to the envy-efficiency trade-off.
2.3 Discussion of Modeling Assumptions
We conclude the section with a discussion of our modeling assumptions.
Age-dependent quality considerations.
While our model does not explicitly account for freshness considerations (i.e., allocating an item that perishes tomorrow is considered to be the same “good” outcome as allocating an item that perishes in 10 days), one natural way to take such considerations into account would be to define a pseudo-perishing time $\widetilde{T}_{b}=T_{b}-a_{b}$, where $a_{b}>0$ is an item-dependent constant that guarantees a minimum amount of time that an item lasts in the household once it is allocated. (In practice, the decision-maker may want to set $a_{b}$ in a category-specific way, such that the allocated item meets a minimum quality threshold at the time of allocation. For instance, milk can last 5-7 days after its “sell-by” date if stored properly in a refrigerator (U.S. Dairy 2025). In this case, it may be desirable to set $a_{b}=3$ in order to ensure that a household can consume the milk for at least 3 days.) Defining all metrics with respect to these pseudo-perishing times, our results would continue to hold, thereby guaranteeing some notion of freshness in our end allocation. [^1] If $\sigma$ is chosen according to the distribution of the pseudo-perishing time $\widetilde{T}_{b}$, the choice of $a_{b}$ may also have an impact on $\sigma$. For instance, if items are allocated in increasing order of $\mathbb{E}[\widetilde{T}_{b}]$, items for which $a_{b}$ is larger (i.e., those for which there is a quality premium) will be allocated earlier. If the decision-maker allocates items according to the variance of $\widetilde{T}_{b}$, on the other hand, given that $a_{b}$ is a constant, this choice would have no impact on $\sigma$.
Alternatively, one could encode freshness considerations by modeling time-dependent utilities for each item. While this would be an interesting direction to explore theoretically, it is less practically relevant with respect to our motivating application of food banks, given that the quantity of goods allocated — as opposed to an estimate of utility-based social welfare —- remains the most important metric reported by these organizations (Prendergast 2022, Orgut et al. 2016).
On the allocation schedule $\sigma$.
In this work we assume the allocation schedule $\sigma$ is an input to the model, as opposed to a lever available to the decision-maker. While treating $\sigma$ as a decision variable would improve both the envy and inefficiency of our algorithm, the reason for this decoupling is two-fold. From a practical perspective, one can imagine that decision-makers have a preference for using interpretable allocation schedules, such as according to expected perishing time or “best-by” date. In contrast, having an allocation schedule depend on the specifics of the guardrail algorithm would be more challenging to communicate to stakeholders. This would be problematic in instances where spoilage occurred, and a unit with a later “best-by” date or expected perishing time was allocated before the unit that spoiled. Such bad events are harder to justify when $\sigma$ is an output of the decision-maker’s optimization problem. By allowing $\sigma$ to be an input to the model, our framework can instead be used by the decision-maker in practice to test the performance of various interpretable allocation schedules, and heuristically determine the one that yields the strongest performance, as we do in our numerical experiments.
We moreover posit that treating $\sigma$ as a decision variable would be analytically challenging, as well as a computationally intractable task. Namely, as we will see in Section 3, optimizing over $\sigma$ reduces to an optimization over $N!$ orderings, which cannot be done efficiently. We however make some progress toward characterizing “high-quality” allocation schedules by $(i)$ providing sufficient conditions on the relationship between $\sigma$ and the perishing distribution for our algorithm’s spoilage to vanish asymptotically (Proposition 4.19), and $(ii)$ numerically showing that allocating units in increasing order of a lower confidence bound on their perishing time is a robust choice for $\sigma$ (Section 5.4). We leave the derivation of structural results surrounding the optimal allocation schedule (which we conjecture will, in general, be highly distribution-dependent), as well as the design of approximately optimal allocation schedules, as interesting open questions.
Finally, we note that the ranking of units is a sufficient statistic for the model we consider in this paper (as opposed to, say, granular forecasts of their perishing times). To see this, note that given a sequence of allocation quantities, the only question that remains for the decision-maker is which units she should allocate in each period. This reduces to the task of finding a ranking of units that minimizes spoilage, given the allocation quantities. One can easily see this when perishing times are known and deterministic; in this case, given an allocation quantity in each period, knowledge of the actual perishing times provides no additional advantage over the ranking that is induced by these true perishing times. This similarly holds for any forecast of units’ perishing times treated as ground truth.
3 Fundamental Limits of Perishability: Lower Bounds
In the presence of perishable resources, a decision-maker must contend with two obstacles: $(i)$ the aggressiveness of the perishing process, and $(ii)$ errors in the perishing prediction $\sigma$. In this section we formalize the impact of these two challenges. Namely, we identify classes of perishing processes for which there is no hope of achieving the optimal fair allocation, $B/N$, and derive lower bounds on any algorithm’s performance as a function of a salient quantity that captures both the aggressiveness of perishing, as well as the quality of the prediction $\sigma$. In the remainder of the section, we say that an online algorithm is feasible over a sample path if it does not run out of budget. Note that, if an algorithm is infeasible over a sample path, it necessarily achieves $\Delta_{\text{\it EF}}=\Theta(1)$.
3.1 On the Aggressiveness of the Perishing Process
We first argue that the proportional allocation $X^{opt}_{t,\theta}=B/N$ is unachievable unless one places restrictions on the rate at which items perish, even under full information over perishing times and demand.
To see this, consider an instance where all items perish at the end of the first round. There is no hope of achieving low envy in this setting, since there are no items left for arrivals from $t=2$ onwards. The following result establishes that the only perishing time realizations for which $B/N$ is achievable are ones in which the fraction of perished items “lags behind” the proportion of arrivals in each period. We formalize this via the notion of offset-expiry, defined below.
Definition 3.1 (Offset-expiring process)
A perishing process $(T_{b})_{t\in[T]}$ is offset-expiring if:
$\frac{P_{<t}}{B}\leq\frac{N_{<t}}{N}\quad\forall\ t\geq 2.$ (5)
We provide an example below.
Example 3.2 (Continued from Example 2.1)
Consider a setting where $T=B=3$, with $N_{1}=1$, $N_{2}=2$, and $N_{3}=1$ (and $N=4$ as a result). Suppose moreover $|\Theta|=1$, with $w_{\theta}=1$. For this sample path, we have:
$\frac{N_{<t}}{N}=\begin{cases}\frac{1}{4}\quad&\text{for }t=2\\ \frac{3}{4}&\text{for }t=3.\end{cases}$ (6)
Applying Equation 5 to this sample path, a process is offset-expiring if and only if:
$P_{<t}\leq\begin{cases}0\quad&\text{for }t=2\\ 2&\text{for }t=3.\end{cases}$ (7)
In words, for any offset-expiring process in this example, we must have $T_{b}\geq 2$ almost surely for all $b\in\mathcal{B}$, with no more than two units perishing at $t=2$ in the worst case.
Theorem 3.3 establishes that offset-expiry exactly captures the trajectories whereby $B/N$ is a feasible allocation when units are allocated in increasing order of $T_{b}$. We refer to this latter ordering as the hindsight optimal ordering. We defer its proof to Appendix 7.
Theorem 3.3
$X_{t,\theta}=B/N$ for all $t,\theta$ is feasible under the hindsight optimal ordering if and only if the perishing process if offset-expiring.
From a managerial perspective, the offset-expiry condition provides decision-makers with guidance on the selection of perishable goods to stock at the beginning of the horizon. Within the context of food banks, for instance, it highlights that rejecting perishable donations outright (as some food banks do) is too severe a policy; rather, the reasonable rule of thumb that says “don’t accept food if it will spoil faster than the rate at which you allocate it” not only suffices, but is the only correct rule of thumb. It moreover underscores the importance of jointly considering the demand and perishing distributions in this selection. We provide necessary and sufficient conditions for offset-expiry to hold with high probability in Section 4.4.
3.2 Unavoidable Loss Due to Perishing
The previous section established that, even under full information about the perishing process, there exist aggressive perishing distributions for which the proportional allocation $B/N$ is not achievable. The following example illustrates how both the inherent aggressiveness of the exogenous perishing process, as well as the quality of the allocation schedule $\sigma$, play an important role in determining what an algorithm can achieve.
Example 3.4 (Continued from Examples 2.1 and 3.2)
Consider the following perishing distributions:
$T_{\texttt{a}}=\begin{cases}1\quad\text{w.p. }1/2\\ 3\quad\text{w.p. }1/2\end{cases}\quad T_{\texttt{b}}=\begin{cases}2\quad\text{w.p. }1/2\\ 4\quad\text{w.p. }1/2\end{cases}\quad T_{\texttt{c}}=\begin{cases}1\quad\text{w.p. }1/2\\ 2\quad\text{w.p. }1/2\end{cases}.$ (8)
Recall, $N_{1}=1$, $N_{2}=2$, and $N_{3}=1$; moreover, $\sigma(\texttt{a})=2$, $\sigma(\texttt{b})=3$, and $\sigma(\texttt{c})=1$. Note that $\sigma$ orders the units in increasing order of expected perishing time for this instance.
By Example 3.2, this process is not offset-expiring for any sample path in which $T_{\texttt{a}}=1$ or $T_{\texttt{c}}=1$. By Theorem 3.3, then, the proportional allocation of $B/N=3/4$ is only feasible on sample paths for which $T_{\texttt{a}}=3$ and $T_{\texttt{c}}=2$, under the hindsight optimal ordering. Even over such sample paths, however, the possibility remains that $B/N$ is infeasible under $\sigma$. To see this, consider the case where $T_{\texttt{a}}=3$, $T_{\texttt{c}}=2$, and $T_{\texttt{b}}=2$. By allocating $3/4$ in each period according to ordering $\sigma$, $2.25$ units will have been allocated by the end of period $2$ (i.e., the entirety of $\texttt{c},\texttt{a}$, and $0.25$ units of b). However, the remainder of b perishes at the end of period $2$, leaving nothing for the individual who arrives in period $3$.
Motivated by this example, our key insight is that, for any static allocation $X$, there exists a high-probability worst-case loss, denoted by $\overline{\Delta}(X)$, that any algorithm incurs, due to both errors in $\sigma$, as well as the inherent aggressiveness of the perishing process. As a result, rather than having a budget of $B$ items, the algorithm has an effective budget of $B-\overline{\Delta}(X)$ items. Under this effective budget, any feasible static allocation must set $X$ such that $\overline{N}X\leq{B-\overline{\Delta}(X)}$, where $\overline{N}$ is a high-probability upper bound on $N$. Noting that $X=0$ is always a feasible solution to this inequality, a natural choice is to set:
$\underline{X}=\sup\left\{X\ \mid\ X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}\right\},$ (9)
if this supremum is achieved. When $\overline{\Delta}(X)=0$ for all $X$ (e.g., either items don’t perish or the process is offset-expiring and predictions are perfect), $\underline{X}=B/\overline{N}$, and we recover the conservative allocation of the no-perishing setting (Sinclair et al. 2023).
By Equation 9, given the worst-case perishing loss $\overline{\Delta}(X)$ for any allocation $X$, one can compute $\underline{X}$ via line search. Obtaining tight bounds on $\overline{\Delta}(X)$, however, is the challenging piece. To see this, note that for any algorithm, the quantity of goods that perished by the end of the horizon is:
$\sum_{t\in[T]}\sum_{b\in\mathcal{B}_{t}^{alg}}(B_{t}^{alg}(b)-X_{t}^{alg}(b))^{+}\cdot\mathds{1}\{T_{b}=t\},$ (10)
where $B_{t}^{alg}(b)$ is the quantity of item $b$ remaining at the beginning of period $t$, and $X_{t}^{alg}(b)$ is the quantity of item $b$ given out in period $t$. Since $X_{t}^{alg}(b)$ depends on the perishing realizations of previous rounds, computing this quantity requires the ability to simulate sufficiently many replications of the static allocation process under $X$, for all $X\in[0,B/\overline{N}]$, and for each of these replications to compute the number of unallocated goods that perished by the end of the horizon under this allocation, an approach which fails to scale.
To tackle this difficulty, it will be useful for us to consider a “slow” consumption process; at a high level, this process will allow us to bound the latest possible time any remaining unit $b$ could be allocated, with high probability. Under this process, $\underline{N}_{\leq t}$ — a high-probability lower bound on $N_{\leq t}$ — individuals arrive before $t+1$, $\underline{N}_{\leq t}{X}$ items are allocated up to period $t\in[T]$, and no items perish. For $b\in\mathcal{B}$, we let ${\tau}_{b}(1\mid X,\sigma)$ be the period in which $b$ would have been entirely allocated under this slow consumption process. Formally,
${\tau}_{b}(1\mid X,\sigma)=\inf\left\{t\geq 1:\underline{N}_{\leq t}X\geq\sigma(b)\right\}.$ (11)
${\tau}_{b}(1\mid X,\ \sigma)$ represents an upper bound on the time an algorithm using static allocation $X$ would allocate $b$, since items ranked higher than $b$ may have perished, thus decreasing the time at which $b$ is allocated. We provide an example of the ${\tau}_{b}(1\mid X,\sigma)$ construction below.
Example 3.5 (continued from Examples 2.1, 3.2 and 3.4)
Fix $X=3/4$. Under deterministic demand, $\underline{N}_{\leq 1}=1$, $\underline{N}_{\leq 2}=3$, and $\underline{N}_{\leq 3}=4$. Then, the latest possible allocation times are given by:
[…]
In words, under the slow consumption process, the latest possible allocation time for the entirety of unit c is $t=2$, since only $0.75$ units have been consumed in the first round. At $t=2$, $2.25$ units have been consumed, meaning that c and a have both been consumed by the end of the second round. Finally, at $t=3$, 33 units have been consumed, meaning that b must also have been consumed by the end of the third round.
Note that a necessary condition for any unit $b$ to have spoiled under static allocation $X$ is for its perishing time $T_{b}$ to be such that $T_{b}<\min\{T,\tau_{b}(1\mid{X},\sigma)\}$. This follows from the fact that, if $T_{b}$ exceeds the latest possible allocation time $\tau_{b}(1\mid{X},\sigma)$, then $b$ must have been allocated before it perished, by definition. Accordingly, we define $\mu(X)=\sum_{b\in\mathcal{B}}\mathbb{P}\left(T_{b}<\min\left\{T,{\tau}_{b}(1\mid X,\sigma)\right\}\right)$ to be an upper bound on the expected number of units that spoil under $X$, and use this quantity to define the worst-case loss $\overline{\Delta}(X)$. Formally:
$\overline{\Delta}({X})=\min\left\{B,\mu(X)+\textsc{Conf}^{P}_{1}(\mu(X))\right\},$ (12)
where $\textsc{Conf}^{P}_{1}(\mu(X))$ is an appropriately chosen confidence bound, to be specified later.
Figure 2: Illustrating the $\underline{X}$ construction for Example 3.6. The dashed line corresponds to the line $Y=X$, and the solid line to $(B-\overline{\Delta}(X))/\overline{N}$. Here, $\underline{X}$ is represented by the red star, the point at which the solid and dashed lines intersect.
With this machinery in hand, we illustrate the $\underline{X}$ construction for our running example.
Example 3.6 (Continued from Examples 2.1, 3.2, 3.4 and 3.5)
Suppose the perishing distributions are as in Example 3.4, $\overline{N}=4$, and $\textsc{Conf}_{1}^{P}(\mu(X))=0$ for all $X$. Let $X=B/\overline{N}=3/4$. Continuing on from Example 3.5, we have:
$\mu(X)=\mathbb{P}\left(T_{\texttt{a}}<2\right)+\mathbb{P}\left(T_{\texttt{b}}<3\right)+\mathbb{P}\left(T_{\texttt{c}}<2\right)=3/2\implies\overline{\Delta}(X)=3/2.$
Then, $\frac{B-\overline{\Delta}(X)}{\overline{N}}=\frac{3-3/2}{4}=3/8<X$, meaning that setting an allocation of $X=3/4$ is infeasible in this example. In this case, the solution to Equation 9 is $\underline{X}=0.375$, as illustrated in Figure 2.
The above example highlights that perishability forces the decision-maker to allocate less than the asymptotically optimal perishing-free allocation of $B/\overline{N}$, in order to guarantee envy-freeness. The notion of unavoidable perishing-induced loss, which we next define, uses the $\underline{X}$ construction to quantify how much less the decision-maker must allocate to account for spoilage that is out of her control.
Definition 3.7 (Perishing-induced loss)
We define $\mathcal{L}^{\textsf{perish}}=\frac{B}{\overline{N}}-\underline{X}$ to be any algorithm’s perishing-induced loss. We moreover term $T\mathcal{L}^{\textsf{perish}}$ to be the cumulative perishing-induced loss.
Theorem 3.8 establishes that any online algorithm’s performance necessarily scales with $\mathcal{L}^{\textsf{perish}}$.
Theorem 3.8
There exists an instance such that, for any online algorithm that is feasible with probability at least $\alpha$, the following holds with probability at least $\alpha$:
$\Delta_{\text{\it efficiency}}\geq T\mathcal{L}^{\textsf{perish}}\quad\quad\Delta_{\text{\it EF}}\geq\mathcal{L}^{\textsf{perish}}.$
Proof 3.9
Proof. Consider an instance with $B=T$, $|\Theta|=1$ and $N_{t}=1$ for all $t$. Suppose moreover that resources have deterministic perishing times, with $T_{b}=b$ for all $b$, and $\sigma(b)=T+1-b$. Since the perishing and demand processes are deterministic, we let $\textsc{Conf}_{1}^{P}(\mu(X))=0$ for all $X$, and $\overline{N}=N$. For ease of notation, we omit the dependence of all quantities on $\theta$ in the remainder of the proof. The following lemma states that under this flipped ordering, any online algorithm is severely limited in its total allocation.
Lemma 3.10
Any feasible algorithm must have $\sum_{t}X_{t}^{alg}\leq 1$.
Proof 3.11
[…]
Lemma 3.10 implies that any feasible static algorithm, which allocates a fixed amount $X_{t}^{alg}=X$ for all $t$, must have $X\leq\frac{1}{N}$. We use this fact to bound $\overline{\Delta}(X)$, for any feasible static allocation.
Lemma 3.12
For any $0<X\leq\frac{1}{N}$, $\overline{\Delta}(X)\geq T-1$.
Proof 3.13
Proof. By definition:
${\tau}_{b}(1\mid X,\sigma)=\inf\{t>0:N_{\leq t}X\geq\sigma(b)\}=\inf\{t>0:tX\geq\sigma(b)\}=\lceil\frac{\sigma(b)}{X}\rceil=\lceil\frac{T+1-b}{X}\rceil$
$\geq T(T+1-b),$
where the final inequality follows from the fact that $X\leq\frac{1}{N}=\frac{1}{T}$. Hence,
$\overline{\Delta}(X)=\sum_{b}\mathds{1}\left\{T_{b}<\min\{T,{\tau}_{b}(1\mid X,\sigma)\}\right\}\geq\sum_{b}\mathds{1}\left\{b<\min\{T,T({T+1-b})\}\right\}=\sum_{b}\mathds{1}\left\{b<T\right\}$
$=T-1,$
[…]
Putting Lemmas 3.10 and 3.12 together, we have:
$\underline{X}:=\sup\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{N}\}\leq\sup\{X\mid X\leq\frac{T-(T-1)}{T}\}=\frac{1}{T}\\ \implies\mathcal{L}^{\textsf{perish}}=1-1/T,$
where the final equality follows from the fact that $\overline{\Delta}(X)=T-1$ for $X=1/T$, and therefore $X\leq\frac{B-\overline{\Delta}(X)}{N}$ is tight at $X=1/T$.
[…]
We conclude the section by noting the implicit dependence of $\mathcal{L}^{\textsf{perish}}$ on the ordering $\sigma$. While $\sigma$ is an input to our model, Theorem 3.8 implies that, if the decision-maker were to optimize over the allocation ordering, the correct choice would be to set $\sigma^{*}\in\arg\min_{\sigma}\mathcal{L}^{\textsf{perish}}(\sigma),$ where we emphasize the dependence of $\mathcal{L}^{\textsf{perish}}$ on the order $\sigma$. Computing such a $\sigma^{*}$, however, is infeasible given the space of $N!$ orderings, as discussed in Section 2.3. In Section 4.4 we identify sufficient conditions on the perishing process and allocation order that guarantee that $\mathcal{L}^{\textsf{perish}}$ vanish asymptotically. We then use these conditions to identify practical and easily interpretable allocation schedules in our numerical experiments (see Section 5.4).
Remark 3.14
We henceforth assume for simplicity that the supremum on the right-hand side of Equation 9 is attained. This is without loss to our results, since our bounds depend on the perishing-induced loss $\mathcal{L}^{\textsf{perish}}$. Hence, if the supremum fails to be attained, one can define $\underline{X}=X^{*}-\epsilon$, where $X^{*}$ is the point of discontinuity and $\epsilon=o(1)$ guarantees feasibility of $\underline{X}$.
Remark 3.15
Since $N_{\leq t}\geq t$ for all $t\in[T]$ one can similarly define ${\tau}_{b}(1\mid X,\sigma)=\inf\{t\geq 1:tX\geq\sigma(b)\}=\lceil\frac{\sigma(b)}{X}\rceil$ as an upper bound on the latest possible perishing time. This quantity can then be interpreted as the “effective rank” of item $b$. This interpretable simplification comes at the cost of our algorithm’s practical performance, but does not affect our subsequent theoretical bounds.
4 Achieving the Optimal Trade-Off Via Perishing-Aware Guardrails
In this section we design an algorithm that achieves the lower bounds described above. We moreover instantiate our bounds for important special cases.
4.1 Algorithm Description
Our algorithm, Perishing-Guardrail, builds on the guardrail algorithm proposed by Sinclair et al. (2023) for the no-perishing setting. In particular, our algorithm always allocates the same amount to all individuals in a given round (independent of their type). As a result, the only source of envy is due to differences in allocations across rounds, as opposed to across individuals within the same round. Our algorithm controls this source of envy via an input parameter $L_{T}=o(1)$; it then tackles the problem of how much to allocate in each round in three main steps:
  1. 1.
    Constructing a static allocation (also referred to as baseline allocation or lower guardrail), $\underline{X}\in\mathbb{R}_{+}$.
  2. 2.
    Setting an “aggressive” allocation $\overline{X}=\underline{X}+L_{T}$ to improve efficiency.
  3. 3.
    Determining an appropriate threshold condition that indicates when to allocate aggressively (i.e., in which periods to allocate $\overline{X}$ instead of $\underline{X}$).
At a high level, the lower guardrail $\underline{X}$ is constructed to be the largest feasible allocation for which the hindsight envy is zero (i.e., all individuals receive the same amount in each period), with high probability. For this to hold, it must be that the algorithm never runs out of budget under $\underline{X}$. This desideratum is precisely satisfied by the static allocation defined in Equation 9, i.e.,
$\underline{X}=\sup\left\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}\right\},$
where we let $\overline{N}=\mathbb{E}[N]+\textsc{Conf}_{0,T}^{N}$ for an appropriately defined confidence term $\textsc{Conf}_{0,T}^{N}$.
While $\underline{X}$ is budget-respecting with high probability, allocating a single budget-respecting quantity in each period may result in significant losses to efficiency. Herein lies the motivation behind the use of the “aggressive” allocation $\overline{X}=\underline{X}+L_{T}$: in periods in which the algorithm “feels” as though it is holding onto more units than it will need throughout the remainder of the horizon, it allocates aggressively. More specifically, in each period, our algorithm verifies whether there is enough budget remaining to accommodate $(i)$ the aggressive allocation $\overline{X}$ in the current period, $(ii)$ the lower guardrail $\underline{X}$ in all future periods, under high demand, and $(iii)$ aggressive future perishing. If enough budget remains, it allocates $\overline{X}$; otherwise, it defaults to $\underline{X}$. Formally, in any given period $t$, given a pessimistic forecast of future perishing $\overline{P}_{t}$, our threshold condition is as follows:
$B_{t}^{alg}\geq N_{t}\overline{X}+\overline{N}_{>t}\underline{X}+\overline{P}_{t},$
where we let $\overline{N}_{>t}=\mathbb{E}[N_{>t}]+\textsc{Conf}_{t,T}^{N}$ for an appropriately defined confidence term $\textsc{Conf}_{t,T}^{N}$.
Our main challenge lies in the construction of $\overline{P}_{t}$. In particular, computing a tight upper bound on future perishing poses significant challenges that do not exist in the setting without perishable resources. In this latter setting, it is easy to obtain tight upper bounds on future demand since demand uncertainty is exogenous, i.e., it is invariant to the decisions made by the algorithm. On the other hand, uncertainty around future perishing is endogenous: as discussed in Section 3, though the distribution around perishing times is fixed, how many — and which — items perish depends heavily on the rate at which items are being allocated, which itself depends on the rate at which items perish. We sidestep this challenge by turning to the “slow” consumption process introduced in Section 3. Namely, fix a round $t\in[T]$, and let $\tau_{b}(t\mid\underline{X},\sigma)$ denote the latest possible allocation time of unit $b$, starting from time $t$. (This latest possible allocation time is a time-dependent generalization of the latest possible allocation time defined in Section 3.2; we formally define it below.) In line with the intuition provided earlier, a necessary condition for $b$ to have spoiled after $t$ is for its perishing time $T_{b}$ to be such that $t\leq T_{b}<\min\{T,\tau_{b}(t\mid\underline{X},\sigma)\}$. We use this necessary condition to obtain a high-probability pessimistic estimate of the number of units that perish after period $t$, which we denote by $\overline{P}_{t}$. Crucially, $\overline{P}_{t}$ is independent of our algorithm’s decisions, which will allow for its tractable analysis. We formalize the construction of this key quantity below.
For all $t^{\prime}\geq t$, let $\underline{N}_{[t,t^{\prime}]}=\mathbb{E}\left[N_{[t,t^{\prime}]}\right]-\textsc{Conf}_{t,t^{\prime}}^{N}$ be a high-probability lower bound on the number of arrivals between $t$ and $t^{\prime}$, for an appropriately defined confidence term $\textsc{Conf}_{t,t^{\prime}}^{N}$. In line with the discussion in Section 3.2, with high probability, $\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}$ corresponds to the smallest quantity of units that could have been consumed by $t^{\prime}$, since any units perishing before $t^{\prime}$ would only have increased consumption. Consider now any unit $b$ in the set of remaining items at the beginning of period $t$. Under the slow process defined above, if $\underline{N}_{<t}+\underline{N}_{[t,t^{\prime}]}\underline{X}\geq\sigma(b)$, then $b$ must have been allocated by $t^{\prime}\geq t$, according to the allocation schedule $\sigma$. We use this intuition to define $\tau_{b}(t\mid\underline{X},\sigma)$, i.e.,
${\tau}_{b}(t\mid\underline{X},\sigma)=\inf\left\{t^{\prime}\geq t:\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}\geq\sigma(b)\right\}.$ (13)
Let $\overline{\mathcal{B}}_{t}$ denote the set of items remaining in period $t$ under this slow process. Via similar logic, $\overline{\mathcal{B}}_{t}=\{\sigma^{-1}(\lceil\underline{N}_{<t}\underline{X}\rceil),\ldots,\sigma^{-1}(B)\}$. Then, an upper bound on the expected number of items that perish from $t$ onwards is given by:
$\eta_{t}=\sum_{b\in\overline{\mathcal{B}}_{t}}\mathbb{P}\left(t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}\right).$ (14)
Here, we deliberately use $\overline{\mathcal{B}}_{t}$ — as opposed to $\mathcal{B}^{alg}_{t}$, which is endogenous to the allocation process — so that our pessimistic perishing estimate is completely exogenous to our algorithm’s decisions. This will be critical in our analysis, allowing us to sidestep a union bound over exponentially many subsets of remaining items.
With this machinery in hand, we define our high-probability pessimistic forecast $\overline{P}_{t}$ as follows:
$\begin{cases}\overline{P}_{t}&=\min\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})\right\}\quad\forall\ t\in[T],\\ \overline{P}_{0}&=B,\end{cases}$ (15)
for an appropriately defined confidence term $\textsc{Conf}_{t}^{P}(\cdot)$. Note that $\overline{P}_{1}=\overline{\Delta}(\underline{X})$ by construction, a fact we will use in the analysis of our algorithm.
We present our algorithm, Perishing-Guardrail, in Algorithm 1. Recall, for $t\in[T]$, $\mathcal{B}^{alg}_{t}$ is used to denote the set of remaining resources at the beginning of time $t$, and $B_{t}^{alg}$ the quantity of remaining resources at the beginning of the period. Moreover, $\textsc{PUA}_{t}^{alg}$ denotes the quantity of unallocated items that perished at the end of round $t$.
Input: Budget $B=B_{1}^{alg}$, allocation schedule $\sigma$, envy parameter $L_{T}$, arrival confidence terms $(\textsc{Conf}_{t,t^{\prime}}^{N})_{t,t^{\prime}\in\{0,\ldots,T\}}$, perishing confidence terms $(\textsc{Conf}_{t}^{P}(\cdot))_{t\in\{1,\ldots,T\}}$, and perishing inputs $(\eta_{t})_{t\in[T]}$ given by Equation 14
Output: An allocation $X^{alg}\in\mathbb{R}^{T\times|\Theta|}$
Compute $\underline{X}=\sup\left\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}\right\}$ and set $\overline{X}=\underline{X}+L_{T}$.
for $t=1,\ldots,T$ do
   Compute $\overline{P}_{t}=\min\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})\right\}$ // Compute ‘‘worst-case’’ future perishing
 if $B_{t}^{alg}<N_{t}\underline{X}$ then // insufficient budget to allocate lower guardrail
      Set $X_{t,\theta}^{alg}=\frac{B_{t}^{alg}}{N_{t}}$ for each $\theta\in\Theta$. Allocate items $b\in\mathcal{B}^{alg}_{t}$ according to $\sigma$.
 else if $B_{t}^{alg}-N_{t}\overline{X}\geq\underline{X}(\mathbb{E}\left[N_{>t}\right]+\textsc{Conf}_{t,T}^{N})+\overline{P}_{t}$ then// use upper guardrail
      Set $X_{t,\theta}^{alg}=\overline{X}$ for each $\theta\in\Theta$. Allocate items $b\in\mathcal{B}^{alg}_{t}$ according to $\sigma$.
 else // use lower guardrail
      Set $X_{t,\theta}^{alg}=\underline{X}$ for each $\theta\in\Theta$. Allocate items $b\in\mathcal{B}^{alg}_{t}$ according to $\sigma$.
   Update $B_{t+1}^{alg}=B_{t}^{alg}-N_{t}X_{t}^{alg}-\textsc{PUA}^{alg}_{t}$
end for
return $X^{alg}$
ALGORITHM 1 Perishing-Guardrail
4.2 Performance Guarantee
Theorem 4.1 below provides upper bounds on counterfactual envy and inefficiency, given hindsight envy parameter $L_{T}$.
Theorem 4.1
For $t^{\prime}>t$, define the confidence terms:
  • $\textsc{Conf}^{N}_{t,t^{\prime}}=\sqrt{2(t^{\prime}-t)|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}$
  • $\textsc{Conf}^{P}_{t}(\eta_{t})=\frac{1}{2}\left(\log(3t\log T/\delta)+\sqrt{\log^{2}(3t\log T/\delta)+8\eta_{t}\log(3t\log T/\delta)}\right)$
Then, with probability at least $1-2\delta$, Algorithm 1 achieves:
$\Delta_{\text{\it EF}}\lesssim\max\{L_{T},\mathcal{L}^{\textsf{perish}}+1/\sqrt{T}\}$ $\Delta_{\text{\it efficiency}}\lesssim\min\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}\right\}+T\mathcal{L}^{\textsf{perish}}$
$\textsc{Envy}\lesssim L_{T}$
where $\lesssim$ drops poly-logarithmic factors of $T$, $\log(1/\delta)$, and absolute constants.
In Section 8.1, we show that these bounds are tight relative to the lower bounds in Theorems 2.4 and 3.8. We dedicate the remainder of this section to further parsing the bounds on counterfactual envy and efficiency given the scalings of $L_{T}$ and $\mathcal{L}^{\textsf{perish}}$. We prove Theorem 4.1 in Section 4.3, and instantiate $\mathcal{L}^{\textsf{perish}}$ for important special cases in Section 4.4.
$L_{T}$$\mathcal{L}^{\textsf{perish}}$$T^{-1/2}$$T^{-1/2}$High-Envy,Low-Perishing(Corollary 4.4)Low-Envy(Corollary 4.2)High-Envy,High-Perishing(Corollary 4.3)
Figure 3: Graphical representation of the regimes described in Corollaries 4.2, 4.4 and 4.3. The curve $T\mathcal{L}^{\textsf{perish}}L_{T}=1$, which partitions the two cases of Corollary 4.4, is shown in gray.
The following corollary simplifies our bounds in the “low-envy” setting where $L_{T}\lesssim 1/\sqrt{T}$.
Corollary 4.2 (Low-Envy)
Suppose $L_{T}\lesssim 1/\sqrt{T}$. Then, Perishing-Guardrail achieves with probability at least $1-2\delta$:
$\Delta_{\text{\it EF}}\lesssim\mathcal{L}^{\textsf{perish}}+1/\sqrt{T}$ $\Delta_{\text{\it efficiency}}\lesssim\sqrt{T}+T\mathcal{L}^{\textsf{perish}}.$
Corollary 4.2 implies that there is no efficiency benefit to increasing $L_{T}$ as long as $L_{T}\lesssim 1/\sqrt{T}$. When $\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$, our algorithm incurs $\widetilde{O}(\sqrt{T})$ envy and inefficiency. In this case, these quantities are driven by the exogenous uncertainty in demand. When $\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$, on the other hand, envy and inefficiency are driven by the unavoidable perishing-induced loss.
Corollary 4.3 next considers the “high-envy, high-perishing” setting, on the other extreme of the spectrum.
Corollary 4.3 (High-Envy, High-Perishing)
Suppose $L_{T}\gtrsim 1/\sqrt{T}$ and $\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$. Then, Perishing-Guardrail achieves with probability at least $1-2\delta$:
$\Delta_{\text{\it EF}}\lesssim\max\{L_{T},\mathcal{L}^{\textsf{perish}}\}$ $\Delta_{\text{\it efficiency}}\lesssim T\mathcal{L}^{\textsf{perish}}.$
Similarly in this regime, increasing $L_{T}$ doesn’t guarantee arbitrary gains in efficiency; thus, setting $L_{T}\sim\mathcal{L}^{\textsf{perish}}$ is optimal. We conclude by exploring the more nuanced “high-envy, low-perishing” regime. In this setting, our algorithm’s guarantees depend on whether the efficiency gain from increasing envy, $L_{T}^{-1}$, exceeds $T\mathcal{L}^{\textsf{perish}}$. We defer its proof to Appendix 8.2.1.
Corollary 4.4 (High-Envy, Low-Perishing)
Suppose $L_{T}\gtrsim 1/\sqrt{T}$, $\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$, and $T\mathcal{L}^{\textsf{perish}}\lesssim L_{T}^{-1}$. Then, Perishing-Guardrail achieves with probability at least $1-2\delta$:
$\Delta_{\text{\it EF}}\lesssim L_{T}$ $\Delta_{\text{\it efficiency}}\lesssim L_{T}^{-1}.$
On the other hand, if $T\mathcal{L}^{\textsf{perish}}\gtrsim L_{T}^{-1}$, Perishing-Guardrail achieves with probability at least $1-2\delta$:
$\Delta_{\text{\it EF}}\lesssim L_{T}$ $\Delta_{\text{\it efficiency}}\lesssim T\mathcal{L}^{\textsf{perish}}.$
Thus, in “high-envy, low-perishing” regimes, if the cumulative perishing-induced loss is order-wise dominated by the efficiency gains from envy (otherwise phrased, our allocation schedule $\sigma$ is high-enough quality and the perishing rate is mild), increasing $L_{T}$ allows Perishing-Guardrail to achieve inversely proportional gains in efficiency, as in the setting without perishable resources. One can do this until moving into the regime where $L_{T}^{-1}\lesssim T\mathcal{L}^{\textsf{perish}}$ (i.e., the cumulative perishing-induced loss dominates efficiency gains from envy). At this point, further increasing $L_{T}$ hurts envy, and has no order-wise impact on efficiency. We summarize these four regimes pictorially in Figure 3.
We conclude the section by noting that our result is a strict generalization of Sinclair et al. (2023): in the simplest setting where $\mathcal{L}^{\textsf{perish}}=0$ (i.e., no perishing, or the process is offset-expiring with perfect predictions), we recover the trade-off described in Theorem 2.4.
4.3 Analysis
The proof of Theorem 4.1 is based on three main building blocks:
  • Defining and bounding the “good event” (Section 4.3.1): We first show that, with probability at least $1-2\delta$, the realizations of future arrivals and perishing are no worse than our algorithm’s pessimistic predictions. As a result, it suffices to condition over such “good” sample paths.
  • Establishing feasibility of $\underline{X}$ (Section 4.3.2): For this good event, we show that the static allocation $\underline{X}$ computed at the start of the horizon will never exhaust the budget, despite incurring perishing-induced loss.
  • Improving efficiency via $\overline{X}$ (Section 4.3.3): We next show that the threshold condition guarantees that the algorithm allocates aggressively enough throughout the horizon to ensure high efficiency.
We use these building blocks for our final bounds on envy and efficiency in Section 4.3.4.
4.3.1 Defining and bounding the “good event.”
We analyze the performance of our algorithm under a so-called “good event” $\mathcal{E}$, the intersection of the following two events:
  1. 1.
    $\mathcal{E}_{N}=\left\{|N_{(t,t^{\prime}]}-\mathbb{E}\left[N_{(t,t^{\prime}]}\right]|\leq\textsc{Conf}^{N}_{t,t^{\prime}}\ \forall\ t,t^{\prime}>t\right\},$
  2. 2.
    $\mathcal{E}_{\overline{P}}=\left\{\overline{P}_{t}\geq\textsc{PUA}^{alg}_{\geq t}\,\forall\ t\in\{0,1,\ldots,T\}\right\}$, where $\textsc{PUA}^{alg}_{\geq t}$ denotes the quantity of unallocated items that perished between the end of round $t$ and the end of round T1T-1.
$\mathcal{E}$ represents the event that the arrival process falls close to its mean, and that $\overline{P}_{t}$ is a pessimistic estimate of the unallocated goods that perish in the future.
The following lemma provides the high-probability bound on $\mathbb{P}(\mathcal{E}_{N})$. We defer its proof — which follows from a standard application of Hoeffding’s inequality — to Appendix 8.3.1.
Lemma 4.5
$\mathcal{E}_{N}$ holds with probability at least $1-\delta$.
The main challenge in analyzing $\mathcal{E}$ lies in showing that $\overline{P}_{t}$ is indeed a pessimistic estimate of our algorithm’s future perishing. To upper bound the amount of unallocated resources that perished between $t$ and T1T-1, we must account for both the uncertainty in arrivals and the realized order in which resources perished, and relate these two sources of uncertainty to the time at which the algorithm intended to allocate these resources. Establishing this upper bound hinges upon the careful construction of the “slow” consumption process. In particular, leveraging $\tau_{b}(t\mid\underline{X},\sigma)$ as an upper bound on the time that each unit is allocated by our algorithm, our bound on the number of units that perish from $t$ onwards is a sum of independent random variables that is decoupled from the algorithm’s past (and future) decisions, thereby allowing us to tractably obtain tight bounds on the true perishing process. We formalize these ideas in Lemma 4.6.
Lemma 4.6
Given $\mathcal{E}_{N}$, $\mathcal{E}_{\overline{P}}$ holds with probability at least $1-\delta$.
Proof 4.7
Proof of Lemma 4.6. We prove the claim by induction.
Base case: $t=0$. Since $\overline{P}_{0}=B$ by definition, we have $\overline{P}_{0}\geq\textsc{PUA}^{alg}_{\geq 0}$ trivially.
Inductive step: $t-1\rightarrow t.$ Let $\mathcal{A}_{\tau}$ be the set of items allocated in round $\tau$, for $\tau\geq t$. Since $\textsc{PUA}^{alg}_{\geq t}$ represents the amount of unallocated goods that perished between $t$ and T1T-1, we have:
$\textsc{PUA}^{alg}_{\geq t}{=}\sum_{\tau=t}^{T-1}\sum_{b\in\mathcal{B}^{alg}_{\tau}}\mathds{1}\left\{T_{b}=\tau,b\not\in\mathcal{A}_{\tau}\right\}$ $=\sum_{b\in\mathcal{B}^{alg}_{t}}\mathds{1}\left\{T_{b}\geq t,b\text{ not allocated before }T_{b}\right\}$
$\leq\sum_{b\in\mathcal{B}^{alg}_{t}}\mathds{1}\left\{t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}\right\},$ (16)
where the inequality follows from the fact that $t\leq T_{b}<\min\{T,\tau_{b}(t\mid\underline{X},\sigma)\}$} is a necessary condition for item $b$ to have spoiled.
A similar argument gives us that $\mathcal{B}^{alg}_{t}\subseteq\overline{\mathcal{B}}_{t}$. Plugging this into (16):
$\textsc{PUA}^{alg}_{\geq t}$ $\leq\sum_{b\in\overline{\mathcal{B}}_{t}}\mathds{1}\left\{t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}\right\}.$ (17)
Recall, $\eta_{t}=\sum_{b\in\overline{\mathcal{B}}_{t}}\mathbb{P}(t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\})$. Applying a Chernoff bound (see Corollary 10.3) to the right-hand side of (17), we obtain that, with probability at least $1-\delta/(3t\log(T))$:
$\textsc{PUA}^{alg}_{\geq t}$ $\leq\eta_{t}+\frac{1}{2}\left(\log(3t\log(T)/\delta)+\sqrt{\log^{2}(3t\log(T)/\delta)+8\eta_{t}\log(3t\log(T)/\delta)}\right)$
$=\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t}).$
Moreover, $\textsc{PUA}^{alg}_{\geq t}\leq\textsc{PUA}^{alg}_{\geq t-1}\leq\overline{P}_{t-1}$, where the second inequality follows from the inductive hypothesis. Putting these two facts together, we obtain:
$\textsc{PUA}^{alg}_{\geq t}\leq\min\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})\right\}=\overline{P}_{t}$
[…]
Lemma 4.8 follows from these high-probability bounds. We defer its proof to Appendix 8.3.2.
Lemma 4.8
Let $\mathcal{E}=\mathcal{E}_{N}\cap\mathcal{E}_{\overline{P}}$. Then, $\mathbb{P}(\mathcal{E})\geq 1-2\delta$.
In the remainder of the proof, it suffices to restrict our attention to $\mathcal{E}$.
4.3.2 Feasibility of $\underline{X}$.
We now show that, given $\mathcal{E}$, our algorithm never runs out of budget, and as a result always allocates $X_{t,\theta}^{alg}\in\{\underline{X},\overline{X}\}$. Since $X_{t,\theta}^{alg}=X_{t,\theta^{\prime}}^{alg}$ for all $\theta,\theta^{\prime}$, for ease of notation in the remainder of the proof we omit the dependence of $X_{t,\theta}^{alg}$ on $\theta$.
Lemma 4.9
Under event $\mathcal{E}$, $B_{t}^{alg}\geq N_{\geq t}\underline{X}$ for all $t\in[T]$.
Proof 4.10
Proof of Lemma 4.9. By induction on $t$.
Base Case: $t=1$. By definition:
$\underline{X}\leq\frac{B-\overline{\Delta}(\underline{X})}{\overline{N}}\implies B\geq\overline{N}\underline{X}+\overline{\Delta}(\underline{X})\geq N\underline{X},$
where the final inequality follows from $\overline{N}\geq N$ under $\mathcal{E}$, and $\overline{\Delta}(\underline{X})\geq 0$.
Step Case: $t-1\rightarrow t$. We condition our analysis on $(X_{\tau}^{alg})_{\tau<t}$, the algorithm’s previous allocations.
Case 1: $X_{\tau}^{alg}=\underline{X}$ for all $\tau<t$. By the recursive budget update, $B_{t}^{alg}=B-N_{<t}\underline{X}-\textsc{PUA}^{alg}_{<t}$, where $\textsc{PUA}^{alg}_{<t}$ denotes the quantity of unallocated goods that perished before the end of round $t$. To show that $B_{t}^{alg}\geq N_{\geq t}\underline{X}$, it then suffices to show that $B-\textsc{PUA}^{alg}_{<t}\geq N\underline{X}$. We have:
$\textsc{PUA}^{alg}_{<t}\leq\textsc{PUA}^{alg}_{\geq 1}\leq\overline{P}_{1}=\overline{\Delta}(\underline{X}),$
where the final inequality follows from Lemma 4.6. Under $\mathcal{E}$, then, as in the base case:
$B-\textsc{PUA}^{alg}_{<t}\geq B-\overline{\Delta}(\underline{X})\geq N\underline{X}.$
Case 2: There exists $\tau<t$ such that $X_{\tau}^{alg}=\overline{X}$. Let $t^{*}=\sup\{\tau<t:X_{\tau}^{alg}=\overline{X}\}$ be the most recent time the algorithm allocated $\overline{X}$. Again, by the recursive budget update:
$B_{t}^{alg}=B_{t^{*}}^{alg}-N_{t^{*}}\overline{X}-N_{(t^{*},t)}\underline{X}-\textsc{PUA}^{alg}_{[t^{*},t)}.$
Since $\overline{X}$ was allocated at $t^{*}$, it must have been that $B_{t^{*}}^{alg}\geq N_{t^{*}}\overline{X}+\overline{N}_{>t^{*}}\underline{X}+\overline{P}_{t^{*}}$. Plugging this into the above and simplifying:
$B_{t}^{alg}$ $\geq N_{t^{*}}\overline{X}+\overline{N}_{>t^{*}}\underline{X}+\overline{P}_{t^{*}}-N_{t^{*}}\overline{X}-N_{(t^{*},t)}\underline{X}-\textsc{PUA}^{alg}_{[t^{*},t)}$
$=\overline{N}_{>t^{*}}\underline{X}-N_{(t^{*},t)}\underline{X}+\overline{P}_{t^{*}}-\textsc{PUA}^{alg}_{[t^{*},t)}$
$\geq N_{\geq t}\underline{X}+\overline{P}_{t^{*}}-\textsc{PUA}^{alg}_{[t^{*},t)},$
[…]
4.3.3 Improving efficiency via $\overline{X}$.
Having established that the algorithm never runs out of budget, it remains to investigate the gains from allocating $\overline{X}$. By the threshold condition, whenever the algorithm allocates $\overline{X}$ it must be that there is enough budget remaining to allocate $\overline{X}$ in the current period, and $\underline{X}$ in all future periods, under high demand and high perishing. Thus, at a high level, $\overline{X}$ being allocated is an indication that the algorithm has been inefficient up until round $t$. The following lemma provides a lower bound on the last time the algorithm allocates $\underline{X}$. This lower bound will later on allow us to establish that, for most of the time horizon, the remaining budget is low relative to future demand, ensuring high efficiency.
Lemma 4.11
Given $\mathcal{E}$, let $t_{0}=\sup\{t:X_{t}^{alg}=\underline{X}\}$ be the last time that $X_{t}^{alg}=\underline{X}$ (or else 0 if the algorithm always allocates according to $\overline{X}$). Then, for some $\tilde{c}=\widetilde{\Theta}(1)$,
$t_{0}>T-\tilde{c}\left(\frac{1}{L_{T}}+\sqrt{\frac{T\mathcal{L}^{\textsf{perish}}}{L_{T}}}\right)^{2}.$
We defer the proof of Lemma 4.11 to Appendix 8.3.3. Observe that, as $\mathcal{L}^{\textsf{perish}}$ increases, our algorithm stops allocating $\underline{X}$ earlier on. We will next see how this loss propagates to our final efficiency bound.
4.3.4 Putting it all together.
With these building blocks in hand, we prove our main result.
Proof 4.12
Proof of Theorem 4.1. By Lemma 4.9, the algorithm never runs out of budget under event $\mathcal{E}$, which occurs with probability at least $1-2\delta$. As a result $X_{t,\theta}^{alg}\in\{\underline{X},\overline{X}\}$ for all $t\in[T]$, $\theta\in\Theta$. We use this to bound envy and efficiency.
Counterfactual Envy: Recall, $\Delta_{\text{\it EF}}=\max_{t,\theta}|w_{\theta}(X_{t,\theta}^{alg}-\frac{B}{N})|\leq w_{\max}\cdot|X_{t,\theta}-\frac{B}{N}|$. We consider two cases.
Case 1: $\underline{X}\leq\overline{X}\leq\frac{B}{N}$. By definition:
$\frac{B}{N}-\underline{X}=\frac{B}{N}-\frac{B}{\overline{N}}+\frac{B}{\overline{N}}-\underline{X}\leq\frac{B}{\underline{N}}-\frac{B}{\overline{N}}+\mathcal{L}^{\textsf{perish}},$
where the inequality follows from the fact that $\underline{N}\leq N$ under $\mathcal{E}$, and $\mathcal{L}^{\textsf{perish}}=B/\overline{N}-\underline{X}$ by definition. We turn our attention to the first two terms:
$\frac{B}{\underline{N}}-\frac{B}{\overline{N}}$ ${=}\frac{B}{\mathbb{E}\left[N\right]-\textsc{Conf}_{0,T}^{N}}-\frac{B}{\mathbb{E}\left[N\right]+\textsc{Conf}_{0,T}^{N}}$
$=\frac{B}{\mathbb{E}\left[N\right]}\left(\frac{1}{1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}-\frac{1}{1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}\right)=\beta_{avg}\left(\frac{1}{1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}-\frac{1}{1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}}\right).$
Using the fact that $\textsc{Conf}_{0,T}^{N}=\sqrt{2T|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}$ and $\mathbb{E}\left[N\right]=\Theta(T)$, there exists $c_{1},c_{2}=\widetilde{\Theta}(1)$ such that, for large enough $T$, $\left(1-\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}\right)^{-1}\leq\left(1-c_{1}/\sqrt{T}\right)^{-1}\leq 1+2c_{1}/\sqrt{T}$ and $\left(1+\frac{\textsc{Conf}_{0,T}^{N}}{\mathbb{E}\left[N\right]}\right)^{-1}\geq\left(1+c_{2}/\sqrt{T}\right)^{-1}\geq 1-c_{2}/\sqrt{T}$. Plugging this into the above:
$\frac{B}{\underline{N}}-\frac{B}{\overline{N}}$ $\leq\beta_{avg}\left(1+2c_{1}/\sqrt{T}-(1-c_{2}/\sqrt{T})\right)\leq\beta_{avg}(2c_{1}+c_{2})/\sqrt{T}\lesssim 1/\sqrt{T}.$
Thus, we obtain $|X_{t,\theta}^{alg}-B/N|\lesssim 1/\sqrt{T}+\mathcal{L}^{\textsf{perish}}$.
Case 2: $\underline{X}\leq\frac{B}{N}\leq\overline{X}$. We have: $|X_{t,\theta}^{alg}-B/N|=\max\left\{\frac{B}{N}-\underline{X},\overline{X}-\frac{B}{N}\right\}\leq\overline{X}-\underline{X}=L_{T}.$
Combining these two cases, we obtain $\Delta_{\text{\it EF}}\lesssim\max\{1/\sqrt{T}+\mathcal{L}^{\textsf{perish}},L_{T}\}$.
Hindsight Envy: Envy is trivially bounded above by $w_{max}\cdot L_{T}\lesssim L_{T}$ since, for any $t,t^{\prime}$:
$w_{\theta}(X_{t^{\prime},\theta}-X_{t,\theta})\leq w_{max}(\overline{X}-\underline{X})=w_{max}L_{T}.$
Efficiency: Let $t_{0}=\sup\{t:X_{t}^{alg}=\underline{X}\}$. Then:
$\Delta_{\text{\it efficiency}}$ $=B-\sum_{t,\theta}N_{t,\theta}X_{t,\theta}=B-\sum_{t}N_{t}X_{t}^{alg}$
$=B_{t_{0}}+\sum_{t<t_{0}}N_{t}X_{t}^{alg}+\textsc{PUA}^{alg}_{<t_{0}}-\sum_{t}N_{t}X_{t}^{alg}$
$=B_{t_{0}}-\sum_{t\geq t_{0}}N_{t}X_{t}^{alg}+\textsc{PUA}^{alg}_{<t_{0}}$
$<N_{t_{0}}\overline{X}+\overline{N}_{>t_{0}}\underline{X}+\overline{P}_{t_{0}}-N_{t_{0}}\underline{X}-N_{>t_{0}}\overline{X}+\textsc{PUA}^{alg}_{<t_{0}}$
$=\underline{X}(\overline{N}_{>t_{0}}-N_{>t_{0}})-(\overline{X}-\underline{X})(N_{>t_{0}}-N_{t_{0}})+\overline{P}_{t_{0}}+\textsc{PUA}^{alg}_{<t_{0}},$
where the inequality follows from $X_{t_{0}}^{alg}=\underline{X}$ and the threshold condition for allocating $\overline{X}$.
Noting that $\underline{X}\leq\beta_{avg}$ and $\overline{N}_{>t_{0}}-N_{>t_{0}}\leq 2\textsc{Conf}_{t_{0},T}^{N}$, we have:
$\underline{X}(\overline{N}_{>t_{0}}-N_{>t_{0}})$ $\leq\beta_{avg}\cdot 2\sqrt{2({T}-{t_{0}})|\Theta|\rho_{max}^{2}\log(2T^{2}/\delta)}$
$\leq 2\beta_{avg}\sqrt{2\tilde{c}|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}\min\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}\right\},$ (18)
where the second inequality follows from Lemma 4.11.
We loosely upper bound the second term by:
$-(\overline{X}-\underline{X})(N_{>t_{0}}-N_{t_{0}})\leq(\overline{X}-\underline{X})N_{t_{0}}\leq L_{T}|\Theta|(\mu_{max}+\rho_{\max}).$ (19)
Finally, consider $\overline{P}_{t_{0}}+\textsc{PUA}_{<t_{0}}^{alg}$. By construction, $\overline{P}_{t_{0}}\leq\overline{P}_{1}=\overline{\Delta}(\underline{X})$. To upper bound $\textsc{PUA}_{<t_{0}}^{alg}$, we consider the process that allocates $\underline{X}$ in each period to all arrivals. Let $B_{t}(\underline{X})$ denote the quantity of remaining items under this process, and $\mathcal{B}_{t}(\underline{X})$ the set of remaining items. We use $\textsc{PUA}_{t}(\underline{X})$ to denote the quantity of unallocated items that perish at the end of period $t$ under this process, and $\textsc{PUA}_{<t}(\underline{X})$ those that perished before the end of period $t$. The following lemma allows us to tractably bound $\textsc{PUA}_{<t_{0}}^{alg}$ via this process. We defer its proof to Appendix 8.3.4.
Lemma 4.13
For all $t\in[T]$,
  1. 1.
    $\mathcal{B}^{alg}_{t}\subseteq\mathcal{B}_{t}(\underline{X})$
  2. 2.
    $\textsc{PUA}_{t}^{alg}\leq\textsc{PUA}_{t}(\underline{X})$.
Using these two facts, we have: $\textsc{PUA}_{<t_{0}}^{alg}\leq\textsc{PUA}_{<t_{0}}(\underline{X})\leq\textsc{PUA}_{\geq 1}(\underline{X})\leq\overline{P}_{1}=\overline{\Delta}(\underline{X}).$ Hence,
$\overline{P}_{t_{0}}+\textsc{PUA}_{<t_{0}}\leq 2\overline{\Delta}(\underline{X})\leq 2\overline{N}\mathcal{L}^{\textsf{perish}}\leq 2\mu_{max}(1+\sqrt{2|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)})T\mathcal{L}^{\textsf{perish}},$ (20)
where the second inequality follows from $\overline{\Delta}(\underline{X})\leq B-\overline{N}\underline{X}=B-\overline{N}\left(\frac{B}{\overline{N}}-\mathcal{L}^{\textsf{perish}}\right)=\overline{N}\mathcal{L}^{\textsf{perish}}$, and the last inequality uses the definition of $\overline{N}$. Putting bounds (4.12), (19) and (20) together, we obtain:
$\Delta_{\text{\it efficiency}}$ $\leq 2\beta_{avg}\sqrt{2\tilde{c}|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)}\min\left\{\sqrt{T},L_{T}^{-1}+\sqrt{TL_{T}^{-1}\mathcal{L}^{\textsf{perish}}}\right\}+L_{T}|\Theta|(\mu_{max}+\rho_{\max})$
$\qquad+2\mu_{max}(1+\sqrt{2|\Theta|\rho_{\max}^{2}\log(2T^{2}/\delta)})T\mathcal{L}^{\textsf{perish}}.$
[…]
Remark 4.14
While Perishing-Guardrail achieves the optimal guarantees for arbitrary $\sigma$ and perishing distributions, the perishing-agnostic guardrail algorithm of Sinclair et al. (2023) is optimal in a special case of our setting. Namely, suppose the decision-maker knows the hindsight optimal ordering, and that there exist deterministic constants $\overline{N}$ and $(\overline{P}_{<t},\underline{N}_{<t},t\geq 2)$, such that
$\mathbb{P}\left(P_{<t}\leq\overline{P}_{<t},N_{<t}\geq\underline{N}_{<t},N\leq\overline{N}\ \forall\ t\geq 2\right)\geq 1-\delta\quad\text{and}\quad\frac{\overline{P}_{<t}}{B}\leq\frac{\underline{N}_{<t}}{\overline{N}}\quad\forall\ t\ \geq 2,$ (21)
for $\delta\in(0,1)$. Under the above condition, offset-expiry holds with probability at least $1-\delta$. Hence, $\mathcal{L}^{\textsf{perish}}=0$ over the sample paths for which (21) holds. As a result, the baseline allocation $\underline{X}=\frac{B}{\overline{N}}$ is feasible over these sample paths. Since allocating aggressively relative to $\underline{X}$ cannot increase the amount of goods that spoil, it follows that the algorithm of Sinclair et al. (2023) is optimal in this setting. While the assumption that the hindsight optimal ordering is known is unlikely to hold in practical settings where the perishing times are stochastic, this special case further highlights the two main features that render this setting challenging: $(i)$ the inherent aggressiveness of the perishing process, and $(ii)$ the unknown perishing order.
4.4 Special Cases
We conclude the section by instantiating our bounds for important special cases. Recall, Theorem 3.3 established that the proportional allocation $B/N$ is achievable if and only if the process is offset-expiring. As a result, for general, non-offset-expiring perishing processes, one should expect $\mathcal{L}^{\textsf{perish}}=\Omega(1)$. In this case, Corollaries 4.2 and 4.3 imply that counterfactual envy is constant, and inefficiency grows linearly in $T$. In this section we restrict our attention to a special class of distributions, parameterized by $\delta\in(0,1)$, which we term $\delta$-offset-expiring, for which $\mathcal{L}^{\textsf{perish}}$ vanishes with respect to $T$ under the hindsight optimal ordering. We formally define this class of processes below.
Definition 4.15 ($\delta$-offset-expiry.)
Fix $\delta\in(0,1)$. A process is $\delta$-offset-expiring if:
${\mathbb{P}\left(\frac{P_{<t}}{B}\leq\frac{N_{<t}}{N}\ \forall\,t\geq 2\right)\geq 1-\delta.}$ (22)
For simplicity, in this section we assume $B=N=T$, with $N_{t}=1$ for all $t$, almost surely. [^2] At the cost of cumbersome algebra, one can relax this assumption and derive entirely analogous results. In this case, Equation 22 reduces to the condition that $\mathbb{P}(P_{<t}\leq t-1\ \forall\ t\geq 2)\geq 1-\delta$.
4.4.1 The i.i.d. setting.
We introduce some additional notation. For $t\in[T]$, we let $\mathcal{B}^{rand}_{<t}=\{b:\mathbb{P}\left(T_{b}<t\right)\in(0,1)\}$ and $\mathcal{B}^{det}_{<t}=\{b:\mathbb{P}\left(T_{b}<t\right)=1\}$. In words, $\mathcal{B}^{det}_{<t}$ represents the set of items that perish before period $t$, almost surely.
Propositions 4.16 and 4.17 respectively establish necessary and sufficient conditions for $\delta$-offset-expiry to hold when the perishing distribution is i.i.d., for non-trivial values of $\delta$.
Proposition 4.16
Suppose there exists $t\geq 2$ such that $\mathbb{E}[P_{<t}]>t-1$. If $\mathcal{B}^{rand}_{<t}=\emptyset$, $\delta$-offset-expiry cannot be satisfied for any value of $\delta\in(0,1)$. Else, $\delta$-offset-expiry cannot be satisfied for $\delta<\frac{1}{2}-{\mathrm{Std}\left[P_{<t}\right]^{-3}}\cdot T$.
Proposition 4.17
Suppose $\mathbb{E}[P_{<t}]\leq t-1$ for all $t\geq 2$. Then, the perishing process is $\delta$-offset-expiring for any $\delta\geq\sum_{t=2}^{T}\min\left\{\left(\frac{\mathrm{Std}\left[P_{<t}\right]}{t-\mathbb{E}[P_{<t}]}\right)^{2},\exp\left(-\frac{{2}(t-\mathbb{E}[P_{<t}])^{2}}{|\mathcal{B}^{rand}_{<t}|}\right)\right\}\mathds{1}\{|\mathcal{B}^{rand}_{<t}|>0\}.$
Propositions 4.16 and 4.17 together illustrate how the mean and variability of the perishing distribution jointly impact offset-expiry. In particular, Proposition 4.17 states that either the expected lag between demand and perishing, given by $t-\mathbb{E}[P_{<t}]$, must be large, or the coefficient of variation (CV) with respect to the random lag process $t-P_{<t}$, must be small. This then reduces to a bound on the variance of $P_{<t}$ in settings where perishing closely tracks demand.
Additionally, note that the necessary condition in Proposition 4.16 fails to hold for one of the most standard models of perishing: geometrically distributed perishing with parameter $1/T$ (that is, a constant fraction of items perish in each period). This highlights that one of the most popular models in the literature is, in a sense, far too pessimistic; for this setting, there is no hope of achieving low envy and efficiency with high probability. Proposition 4.18 below shows how the perishing rate under this popular distribution affects offset-expiry and the perishing-induced loss $\mathcal{L}^{\textsf{perish}}$.
Proposition 4.18
Suppose $T_{b}\sim\emph{Geometric}(p)$ for all $b\in\mathcal{B}$, with $p=o(1/T)$. Then, the perishing process is $\delta$-offset-expiring for any $\delta\geq 2\log T\cdot\frac{Tp}{(1-Tp)^{2}}$. Moreover, $\underline{X}\geq 1-3Tp-\frac{\log(3\log(T)/\delta)}{T}$ for any ordering $\sigma$. As a result,
$\mathcal{L}^{\textsf{perish}}\leq 3Tp+\frac{\log(3\log(T)/\delta)}{T}.$
Figure 4: Maximum feasible allocation $\underline{X}$ vs. $\alpha$, for $T_{b}\sim\emph{Geometric}(T^{-(1+\alpha)})$, $B=300$, $T=150$, $N_{t}$ drawn from a truncated normal distribution $\mathcal{N}(2,0.25)$, and $\delta=1/T$. Here, $\underline{X}$ was calculated via line search, with Monte Carlo simulation used to estimate $\Delta(X)$ for each value of $X$. The dashed line represents the “naive” allocation $B/\overline{N}=0.89$ which ignores possible perishing, and the green line is the curve of best fit to $\underline{X}$.
Proposition 4.18 establishes that, in the worst case, $\underline{X}$ decays linearly in the rate $p$ at which goods spoil. This highlights the extent to which the commonly used (and practically pessimistic) geometric model limits both the kinds of perishable goods selected before allocation, as well as the rate at which a decision-maker can allocate goods. Letting $p=T^{-(1+\alpha)}$, $\alpha\in(0,1)$, Proposition 4.18 implies that $\mathcal{L}^{\textsf{perish}}$ is on the order of $T^{-\alpha}$. Alternatively, if a decision-maker wants no more than $T^{-\alpha}$ loss relative to the proportional allocation, Proposition 4.18 provides an upper bound of $T^{-(1+\alpha)}$ on $p$.
We illustrate the scaling of $\underline{X}$ under geometric perishing numerically in Figure 4, for an instance where the decision-maker also faces demand uncertainty. We observe that $\underline{X}$ is concave increasing in $\alpha$, and that the lower bound in Proposition 4.18 provides a good fit for $\underline{X}$, even under demand uncertainty. For $\alpha$ close to 0 (i.e., $p\sim 1/T$), $\underline{X}\approx 0.4$, less than half of the “naive” no-perishing allocation, $B/\overline{N}=0.89$. For $\alpha=1$, $\underline{X}\approx B/\overline{N}$. Note that, for $\alpha>1$, $\underline{X}$ is limited by the confidence bound $\log(3\log T/\delta)/T$ in Proposition 4.18. Plugging the lower bound on $\delta$ into this term, this implies that, even under deterministic demand, $\mathcal{L}^{\textsf{perish}}=\Omega(1/T)$.
4.4.2 The non-i.i.d. setting.
When the perishing distribution is not identical across different units, the quality of the allocation schedule $\sigma$ can have a significant impact on the perishing-induced loss $\mathcal{L}^{\textsf{perish}}$. Proposition 4.19 illustrates how the allocation schedule and perishing distribution come together in determining $\mathcal{L}^{\textsf{perish}}$.
Proposition 4.19
Suppose there exists $\alpha\in(0,1)$ such that:
  1. 1.
    $\mathbb{E}\left[T_{b}\right]>\min\left\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\right\}$
  2. 2.
    $\sum_{b}\frac{\mathrm{Var}\left[T_{b}\right]}{\left(\mathbb{E}\left[T_{b}\right]-\min\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\}\right)^{2}}\leq\frac{1}{2}T^{1-\alpha}$.
Then, for any $\delta\geq 3\log(T)e^{-\frac{1}{8}T^{1-\alpha}}$, the process is $\delta$-offset-expiring, and $\underline{X}\geq 1-T^{-\alpha}$. As a result, $\mathcal{L}^{\textsf{perish}}\leq T^{-\alpha}$.
Proposition 4.19 highlights the two key components that determine the baseline allocation: $(i)$ the lag between the expected perishing time $\mathbb{E}\left[T_{b}\right]$ and the expected allocation time $\min\{T,\lceil\frac{\sigma(b)}{1-T^{-\alpha}}\rceil\}$ (which we colloquially refer to as “room to breathe”), and $(ii)$ the variance of the perishing time. Specifically, Condition 1 implies that, if $\mathbb{E}\left[T_{b}\right]$ is low, it must be that the item is allocated early on in the horizon (i.e., $\sigma(b)$ is low). This encodes the “race against time” intuition that is typically held around perishing. Condition 2 can be viewed as an upper bound on the cumulative adjusted coefficient of variation of the perishing process. High-variance perishing times and smaller “room to breathe” push $\alpha$ down, resulting in a lower maximum feasible allocation rate, and consequently higher perishing-induced loss. In Section 5.4 we build on this intuition to propose heuristic schedules with low $\mathcal{L}^{\textsf{perish}}$.
5 Numerical Experiments
In this section we study the practical performance of Perishing-Guardrail via an extensive set of numerical experiments. We first consider one of the most popular (and aggressive) models of perishing: geometrically distributed perishing times. We explore the dependence of the envy-efficiency trade-off on the perishing rate for this setting, leveraging the empirical trade-off curves to provide guidance on the choice of envy parameter $L_{T}$. We moreover compare the performance of Perishing-Guardrail to its perishing-agnostic counterpart (Sinclair et al. 2023) on both synthetic and real-world datasets (Keskin et al. 2022). Lastly, for the non-i.i.d. setting we numerically study choices of allocation schedules that achieve the two-fold desideratum of being practical for decision-makers and achieving small perishing-induced loss. See https://github.com/seanrsinclair/perishing_fairness for the code database.
5.1 Description of Metrics and Benchmark Policies
We evaluate Perishing-Guardrail using $\delta=\frac{1}{T}$, benchmarking it against three algorithms from the literature:
  • Vanilla-Guardrail (Sinclair et al. (2023)): a guardrail-based algorithm designed for settings without perishable resources;
  • Static-$\underline{X}$: the algorithm that allocates $X_{t,\theta}=\underline{X}$ for all $t,\theta$, until it runs out of resources;
  • Static-$\frac{B}{\overline{N}}$: the algorithm that allocates $X_{t,\theta}=\frac{B}{\overline{N}}$ for all $t,\theta$, until it runs out of resources.
We report the performance of these algorithms along five metrics: $(i)$ expected counterfactual envy $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$, $(ii)$ expected hindsight envy $\mathbb{E}\left[\textsc{Envy}\right]$, $(iii)$ expected inefficiency $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$, $(iv)$ expected spoilage $\mathbb{E}\left[\textsc{Spoilage}\right]$, and lastly $(v)$ stockout probability $\mathbb{E}\left[\textsc{Stockout}\right]$, i.e., the proportion of replications for which the algorithm runs out of resources before the end of the time horizon.
5.2 Geometric Perishing: Synthetic Experiments
We first consider the setting in which each unit perishes independently with probability $p$ in each period, i.e., $T_{b}\sim\text{Geometric}(p)$, for all $b\in\mathcal{B}$. Throughout the section, we assume $|\Theta|=1$. Since perishing times are identically distributed, the allocation order $\sigma$ does not have any impact on the performance of the algorithm; hence, we assume $\sigma$ is the identity ordering. To study the impact of the perishing rate on algorithm performance, we let $p=T^{-(1+\alpha)}$, varying $\alpha\in\{0.1,0.2,0.25,0.3\}$ (see Section 9 for additional values of $\alpha$). We let $T=100$, $B=2T$, and draw demand $N_{t}$ from a truncated normal distribution $\mathcal{N}(2,0.25)$, for $t\in[T]$. All results are averaged over 150 replications.
Empirical trade-off curves.
(a) $T_{b}\sim\emph{Geometric}(T^{-1.1}):\underline{X}=0.70,$ $\mathcal{L}^{\textsf{perish}}=0.4,\mathbb{P}(\mathcal{E}_{OE})=0.89$.
(b) $T_{b}\sim\emph{Geometric}(T^{-1.2}):\underline{X}=0.84,$ $\mathcal{L}^{\textsf{perish}}=0.26,\mathbb{P}(\mathcal{E}_{OE})=0.97$
(c) $T_{b}\sim\emph{Geometric}(T^{-1.25}):\underline{X}=0.89,$ $\mathcal{L}^{\textsf{perish}}=0.21,\mathbb{P}(\mathcal{E}_{OE})=0.99$
(d) $T_{b}\sim\emph{Geometric}(T^{-1.3}):\underline{X}=0.93,$ $\mathcal{L}^{\textsf{perish}}=0.17,\mathbb{P}(\mathcal{E}_{OE})=0.99$
Figure 5: Empirical trade-off between $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$ and $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$. The points on the trade-off curve correspond to increasing values of $L_{T}$, from left to right. Static-$\frac{B}{\overline{N}}$ and Static-$\underline{X}$ respectively correspond to Vanilla-Guardrail and Perishing-Guardrail for $L_{T}=0$.
We begin by numerically investigating the impact of the perishing rate on the envy-efficiency frontier, and use the empirical trade-off curves to provide guidance on how decision-makers should select $L_{T}$. For these instances, we compare Perishing-Guardrail and Vanilla-Guardrail (Sinclair et al. 2023) with $L_{T}=T^{-\beta}$, $\beta\in\{0,0.05,0.1,\ldots,1\}$.
The empirical trade-off curves for each value of $\alpha$ can be found in Figure 5. In each case, we report the probability that a sample path is offset-expiring, denoted by $\mathbb{P}(\mathcal{E}_{OE})$. For $\alpha\in\{0.1,0.2\}$ (Figures 5(a) and 5(b)), Vanilla-Guardrail makes close to no gains in efficiency for any value of $L_{T}$. In these high-perishing settings, then, setting $L_{T}=0$ is optimal (there is no trade-off), in stark contrast to the classic setting without perishability. As $\alpha$ increases (Figures 5(c) and 5(d)), Vanilla-Guardrail attains small gains in efficiency, but plateaus very quickly. This yields the important insight that perishing-agnostic algorithms are not able to leverage unfairness to improve efficiency in the presence of perishable resources. Perishing-Guardrail, on the other hand, sees extremely large gains in efficiency for a very small increase in envy, across all values of $\alpha$. For larger values of $L_{T}$, we observe marginal gains in efficiency; the horizontal asymptote observed across all values of $\alpha$ is precisely the cumulative perishing-induced loss, $T\mathcal{L}^{\textsf{perish}}$ (Theorem 3.8). Moreover, the vertical asymptote across all plots corresponds to the unavoidable loss due to demand uncertainty (Theorem 2.4).
Note that, for small values of $\alpha$, the Perishing-Guardrail empirical trade-off curve lies to the left of that of Vanilla-Guardrail, i.e., it achieves lower counterfactual envy across all values of $L_{T}$. As we will see below, this is due to the fact that Vanilla-Guardrail achieves extremely high stockout rates. This effect is diminished as $\alpha$ increases (i.e., the perishing rate decreases). As this happens, both curves move down and to the left (and closer) as they achieve lower counterfactual envy and inefficiency due to spoilage. When the perishing rate is negligible ($\alpha=0.3$), the empirical trade-off curve of Vanilla-Guardrail is slightly to the left of that of Perishing-Guardrail; this is due to the loss incurred by our modified $\underline{X}$ construction, which always allocates less than $B/\overline{N}$ as a baseline. However, even when perishing is negligible Perishing-Guardrail is slightly more efficient that Vanilla-Guardrail, despite its baseline allocation being lower. This runs counter to the intuition that Vanilla-Guardrail should be more efficient since it has a higher baseline allocation. The reason for this is the difference in the two algorithms’ threshold allocation decisions. Our algorithm, Perishing-Guardrail, allocates $\overline{X}=\underline{X}+L_{T}$ if it forecasts that it has enough budget remaining to allocate $\overline{X}$ in period $t$, and $\underline{X}$ onwards. On the other hand, Vanilla-Guardrail allocates $B/\overline{N}+L_{T}$ if it has enough budget remaining to allocate this high amount in period $t$, and $B/\overline{N}$ in all future periods. Since $B/\overline{N}>\underline{X}$, Vanilla-Guardrail depletes its budget faster than Perishing-Guardrail whenever they both allocate the lower guardrail. Hence, Perishing-Guardrail is able to allocate aggressively more frequently than Vanilla-Guardrail, which results in improved efficiency and decreased spoilage.
These empirical trade-off curves moreover help to provide guidance on the choice of $L_{T}$. In particular, across all experiments, the cusp of the trade-off curve lies at $L_{T}\sim T^{-0.35}$, which is larger than no-perishing cusp of $L_{T}\sim T^{-1/2}$ (Sinclair et al. 2023). This is due to the fact that our baseline allocation is significantly lower to avoid perishing-induced stockouts; hence, in order to recover from this inefficiency, $L_{T}$ must be higher. We use this observation in the following experiments, comparing the performance of Perishing-Guardrail and Vanilla-Guardrail for $L_{T}\sim T^{-0.35}$.
Benchmark comparisons.
Next, we compare the performance of all algorithms along our five metrics as the number of rounds $T$ grows large, for $\alpha\in\{0.1,0.2,0.3\}$. Based on our prior insights, we use $L_{T}=T^{-0.35}$ for both Perishing-Guardrail and Vanilla-Guardrail. Our results can be found in Figure 6.
(a) $T_{b}\sim\emph{Geometric}(T^{-1.1})$
(b) $T_{b}\sim\emph{Geometric}(T^{-1.2})$
(c) $T_{b}\sim\emph{Geometric}(T^{-1.3})$
Figure 6: Algorithm comparison across $\mathbb{E}\left[\Delta_{\text{\it EF}}\right],\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right],\mathbb{E}\left[\textsc{Envy}\right]$, and $\mathbb{E}\left[\textsc{Spoilage}\right]$, for $\alpha\in\{0.1,0.2,0.25,0.3\}$.
We identify three regimes:
  • High Perishing ($\alpha=0.1$, Figure 6(a)): While unfairness (as measured by $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ and $\mathbb{E}\left[\textsc{Envy}\right]$) is decreasing in $T$ under Perishing-Guardrail and Static-$\underline{X}$, Vanilla-Guardrail and Static-$\frac{B}{\overline{N}}$ perform remarkably poorly along these two metrics. This is due to the fact that these latter algorithms fail to account for the unavoidable perishing loss, resulting in an extremely high stockout probability, as illustrated in Table 1, for $T=150$. In contrast, the two perishing-aware algorithms rarely run out of resources. This underscores the importance of modifying the baseline guardrail $\underline{X}$, which was specifically constructed to avoid stockouts due to unavoidable perishing.
    Comparing Static-$\underline{X}$ to Perishing-Guardrail, our results also demonstrate that, in this high-perishing regime, the strategy of cautiously over-allocating by $L_{T}$ comes at a significant reduction in inefficiency $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$, at close to no increase in counterfactual envy $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$.
  • Medium Perishing ($\alpha=0.2$, Figure 6(b)): Though we observe similar trends as when $\alpha=0.1$, all algorithms perform better across the board. Perishing-Guardrail exhibits slightly higher efficiency and lower spoilage than its perishing-agnostic counterpart in this regime, since it satisfies the threshold condition more frequently, as described above. Still, in Table 1 we see that the perishing-agnostic algorithms run out of resources in over 50% of replications.
  • Low Perishing ($\alpha=0.3$, Figure 6(c)): For this smaller perishing rate, Vanilla-Guardrail stocks out significantly less frequently. Putting this together with the fact that $B/\overline{N}>\underline{X}$, this explains the fact that it has lower counterfactual envy than Perishing-Guardrail. However, along all other metrics Perishing-Guardrail improves upon Vanilla-Guardrail. The improvements in efficiency and spoilage are due to the same effects as described above; moreover, our algorithm improves upon Vanilla-Guardrail on $\mathbb{E}\left[\textsc{Envy}\right]$ since it never stocks out.
$\alpha$ $\frac{B}{\overline{N}} Static$-$\underline{X}$ Vanilla-Guardrail Perishing-Guardrail
$0.1$ $99\pm 0.02 0.0$$00\pm 0.0$ $0\pm 0.0 0.11±0$$1\pm 0.06$
$0.2$ $63\pm 0.1 0.0$$00\pm 0.0$ $68\pm 0.09 0.0$$03\pm 0.03$
$0.25$ $17\pm 0.07 0.0$$00\pm 0.0$ $22\pm 0.08 0.0$$00\pm 0.0$
$0.3$ $3\pm 0.04 0.00±0$$\pm 0.0 0.0$$6\pm 0.05 0.0$$00\pm 0.0$
Table 1: Comparison of stockout probabilities, for $T=150$, $T_{b}\sim\text{Geometric}(T^{-(1+\alpha})$. The second number in each cell corresponds to $95\%$ confidence intervals.
Overall, our results highlight the robustness of Perishing-Guardrail to perishability, as it is able to achieve similar if not improved performance compared to Vanilla-Guardrail in settings where there is limited perishing, with vastly superior performance in high-perishing settings.
Algorithm $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ $\mathbb{E}\left[\textsc{Envy}\right]$ $\mathbb{E}\left[\textsc{Spoilage}\right]$ $\mathbb{E}\left[\textsc{Stockout}\right]$ $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$
Static-$\frac{B}{\overline{N}}$ $1.18\pm 0.01$ $1.03\pm 0.0$ $346.4\pm 2.6$ $1.0\pm 0.0$ $444.6\pm 2.7$
Static-$\underline{X}$ $60\pm 0.01 0.0$$0\pm 0.0$ $475.9\pm 2.7$ $0\pm 0.0$ $605.5\pm 3.0$
Vanilla-Guardrail $1.17\pm 0.01$ $1.44\pm 0.0$ $341.4\pm 3.0$ $1.0\pm 0.0$ $343.5\pm 2.9$
Perishing-Guardrail $0.78\pm 0.04$ $0.42\pm 0.05$ $372.2\pm 3.2$ $0.39\pm 0.09$ $372.7\pm 3.2$
Table 2: Performance of the different algorithms (for $L_{T}=T^{-0.35}$) on the “ginger” dataset in Keskin et al. (2022). The second number in each cell corresponds to 95% confidence intervals.
5.3 Geometric Perishing: Real-World Instance
We next investigate the performance of our algorithm using the real-life dataset provided by Keskin et al. (2022), which tracks daily demand, replenishments, and perishing of ginger at a leading supermarket chain in China over $T=365$ days. We assume each unit’s perishing time is geometrically distributed with parameter $p$. Fitting this geometric perishing model to the data, we obtain an estimate of $p=0.00224$. We similarly fit a truncated normal distribution to the demand data, obtaining $\mathcal{N}(3.2,1.85^{2})$. Finally, we set the total budget to $B=365\cdot 3.2=1168$, corresponding to the total expected demand over the horizon. (See Section 9 for additional details on the data calibration.) For these inputs, $B/\overline{N}=0.89$ and $\underline{X}=0.46$. Under these parameters, the offset-expiry condition is only satisfied 65.2% of the time. Given this aggressive perishing, perishing-agnostic algorithms are expected to perform particularly poorly.
In Table 2 we compare the performance of the different algorithms. We observe the following:
  • As conjectured, Static-$\frac{B}{\overline{N}}$ and Vanilla-Guardrail stock out on 100% of replications since they fail to account for endogenous perishing. The high stockout probabilities of Static-$\frac{B}{\overline{N}}$ and Vanilla-Guardrail lead to high unfairness (vis-à-vis hindsight and counterfactual envy), since later arrivals receive allocations of zero. In contrast, Static-$\underline{X}$ never stocks out. Perishing-Guardrail achieves a higher stockout rate of 40%, due to its more aggressive allocations. Still, our algorithm’s counterfactual envy and hindsight envy are over 30% and 70% lower, respectively, than that of Vanilla-Guardrail.
  • Perishing-Guardrail allocates approximately 10% fewer goods than Vanilla-Guardrail. It is notable, however, that it is more efficient than Static-$\frac{B}{\overline{N}}$; this highlights that naively allocating more aggressively need not always generate efficiency gains.
Overall, we see that even when offset-expiry fails to hold over a significant fraction of sample paths, for small losses in efficiency our algorithm makes major gains in fairness relative to perishing-agnostic algorithms. This highlights the robustness of our algorithm to aggressive perishing.
5.4 Non-i.i.d. Perishing
Theorem 4.1 established that the theoretical performance of Perishing-Guardrail improves for orderings $\sigma$ under which $\mathcal{L}^{\textsf{perish}}$ is lower. In practice, however, $\mathcal{L}^{\textsf{perish}}$ has an extremely subtle impact on our algorithm’s efficiency. Namely, while lower $\mathcal{L}^{\textsf{perish}}$ implies a higher $\underline{X}$, this also results in a lower value of $\overline{P}_{t}$ in each round $t$. The net effect on whether the threshold condition is satisfied in each period, then, remains unclear. In this section we study the impact of reasonable heuristic orderings on the practical performance of our algorithm. Motivated by the insights of Proposition 4.19, we compare the following orderings:
  • Increasing Mean ($\sigma^{\textsf{Mean}}$): Increasing order of $\mathbb{E}\left[T_{b}\right]$;
  • Decreasing Coefficient of Variation (CV) ($\sigma^{\textsf{CV}}$): Decreasing order of $\mathrm{Std}\left[T_{b}\right]/\mathbb{E}\left[T_{b}\right]$. For fixed expected perishing time, this schedule allocates high-variance units earlier on. Conversely, for fixed variance, it allocates items according to the Increasing Mean schedule;
  • Increasing Lower Confidence Bound (LCB) ($\sigma^{\textsf{LCB}}$): Increasing order of $\mathbb{E}\left[T_{b}\right]-1.96\mathrm{Std}\left[T_{b}\right]$. This ordering allocates items according to the lower bound of the 95% confidence interval of the normal approximation to its perishing time. This lower bound is expected to be small if either the expected perishing time is small or the variance is large.
We break ties randomly in all cases.
As in Section 5.2, we draw the demands $N_{t}$ from a truncated normal distribution, $\mathcal{N}(2,0.25)$; we moreover let $T=50$, $B=100$, and $L_{T}={T}^{-0.35}$. Finally, we consider two sets of perishing distributions:
  • Instance 1: Front-loaded variability
    $T_{b}=\begin{cases}\text{Uniform}(T/2-b/2,T/2+b/2)&\quad b\leq T\\ T&\quad b>T\end{cases}$ (23)
  • Instance 2: Back-loaded variability
    $T_{b}=\begin{cases}b+1&\quad b\leq T\\ \text{Uniform}(T,b+1)&\quad b>T\end{cases}$ (24)
Tables 3 and 4 show the performance of our algorithm over these instances, for all three allocation schedules.
Order $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]$ $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ $\mathbb{E}\left[\textsc{Envy}\right]$ $\mathbb{E}\left[\textsc{Spoilage}\right]$ $\mathbb{E}\left[\textsc{Stockout}\right]$ $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$
$\sigma^{\textsf{Mean}}$ $0.10\pm 0.004$ $0.36\pm 0.02$ $0.51\pm 0.02$ $5.78\pm 0.4$ $0.04\pm 0.04$ $6.28\pm 0.4$
$\sigma^{\textsf{CV}}$ / $\sigma^{\textsf{LCB}}$ $0.0\pm 0.0$ $0.44\pm 0.02$ $0.48\pm 0.02$ $1.24\pm 0.1$ $0.06\pm 0.04$ $1.79\pm 0.1$
Table 3: Performance of Perishing-Guardrail for $L_{T}=T^{-0.35}$ on Instance 1 (Eq. 23).
Order $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]$ $\mathbb{E}\left[\Delta_{\text{\it EF}}\right]$ $\mathbb{E}\left[\textsc{Envy}\right]$ $\mathbb{E}\left[\textsc{Spoilage}\right]$ $\mathbb{E}\left[\textsc{Stockout}\right]$ $\mathbb{E}\left[\Delta_{\text{\it efficiency}}\right]$
$\sigma^{\textsf{Mean}}$ / $\sigma^{\textsf{LCB}}$ $0.0\pm 0.0$ $0.41\pm 0.02$ $46\pm 0.02 0.0$$0\pm 0.0$ $0.1\pm 0.05$ $0.56\pm 0.07$
$\sigma^{\textsf{CV}}$ $0.47\pm 0.0$ $0.51\pm 0.05$ $0.48\pm 0.02$ $48.3\pm 0.08$ $0.01\pm 0.0$ $48.7\pm 0.07$
Table 4: Performance of Perishing-Guardrail for $L_{T}=T^{-0.35}$ on Instance 2 (Eq. 24).
For the first instance, the Increasing Mean schedule allocates the first $T$ items uniformly at random, ignoring the fact that, for $b\leq T$, as $b$ increases the item is more likely to perish earlier on in the horizon. The Decreasing CV / Increasing LCB schedules, on the other hand, are identical: they allocate the first $T$ resources in decreasing order of $b$, and allocate the remaining uniformly at random. Notably, the Decreasing CV / Increasing LCB order achieves $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]=0$, i.e., $\underline{X}=B/\overline{N}$, as in the no-perishing setting. (Note that $\mathcal{L}^{\textsf{perish}}=0$ implies that this is an optimal ordering.) Since its baseline allocation is higher it results in 78% less spoilage than the Increasing Mean order, and a 71% decrease in inefficiency. However, this order performs slightly worse with respect to counterfactual envy and stockouts: this is again due to the more aggressive allocations.
For the second instance, the Increasing Mean and Increasing LCB schedules are identical: they allocate items lexicographically. The Decreasing CV schedule, on the other hand, allocates the last $T$ items (in increasing order of $b$) before the first $T$ resources, since $\mathrm{Std}\left[T_{b}\right]=0$ for all $b\leq T$. In this setting, the first schedule is optimal with respect to perishing-induced loss, with $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]=0$. This more aggressive allocation results in a 10% stockout rate (versus 1% for the Decreasing CV schedule), but outperforms the Decreasing CV order across all other metrics. This is intuitive as the number of errors in this latter, clearly bad order results in $\mathbb{E}\left[\mathcal{L}^{\textsf{perish}}\right]=0.47$, approximately $50\%$ of the baseline allocation $B/\overline{N}$. The algorithm then incurs both high inefficiency and spoilage.
These results indicate that the Increasing LCB schedule is both a practical and robust candidate allocation order as it hedges against the inherent variability of the perishing process.
6 Conclusion
This paper considers a practically motivated variant of the canonical problem of online fair allocation wherein a decision-maker has a budget of perishable resources to allocate fairly and efficiently over a fixed time horizon. Our main insight is that perishability fundamentally impacts the envy-efficiency trade-off derived for the no-perishing setting: while a decision-maker can arbitrarily sacrifice on envy in favor of efficiency in this latter setting, this is no longer the case when there is uncertainty around items’ perishing times. We derive strong lower bounds to formalize this insight, which are a function of both the quality of the decision-maker’s prediction over perishing times, as well as the inherent aggressiveness of the perishing process. We moreover design an algorithm that achieves these lower bounds; this algorithm relies on the construction of a baseline allocation that accounts for the unavoidable spoilage incurred by any online algorithm. From a technical perspective, the main challenge that the perishing setting presents is that the uncertainty around the quantity of resources that spoil in the future is endogenous, in contrast to the exogenous uncertainty on the number of arrivals in the classical setting. Deriving tight bounds on spoilage (both for our lower bounds as well as in the design of our algorithm) relied on the “slow allocation” construction, which rendered the highly coupled process amenable to tractable analysis. Our numerical experiments also demonstrate our algorithm’s strong performance against state-of-the-art perishing-agnostic benchmarks.
From a practical perspective, our results provide concrete, intuitive guidance as to the composition of perishable donations that a food bank should accept in order to achieve a fair allocation (in the proportional sense). At a high level, the optimal mix of goods should ensure that cumulative perishing always lags behind cumulative demand. Moreover, we uncover the important — and counterintuitive — insight that guaranteeing a fair allocation in the presence of perishable goods may require the decision-maker to allocate significantly less than the static proportional benchmark. Finally, our numerical results suggest that an allocation schedule that ranks items according to high-probability lower bounds on their perishing times is a robust choice that appropriately hedges against the inherent variability of the perishing process. In particular, we observe that this allocation schedule outperforms natural alternatives such as ordering items according to expected perishing time.
In terms of future directions, our model assumes that the decision-maker allocates items according to a fixed allocation schedule. Though our results do not require that the perishing distribution be memoryless, allowing for time-varying / adaptive allocation schedules, though less practical, would improve our algorithm’s performance in non-memoryless settings. This relates back to the question of deriving theoretical insights into the structure of optimal allocation schedules. Finally, though this paper considered exogenous depletion of the budget, a natural practical extension is one wherein $B$ evolves stochastically, accounting for external donations independent of the allocations made by the algorithm.
[…]
The authors thank the area editor, associate editor, and two anonymous referees whose comments and suggestions were tremendously helpful in improving the paper. The authors would also like to thank Connor Lawless for insightful conversations about this work. Part of this work was done while Sean Sinclair and Sid Banerjee were visiting the Simons Institute for the Theory of Computing for the semester on Data-Driven Decision Processes. We gratefully acknowledge funding from the National Science Foundation under grants ECCS-1847393, DMS1839346, CCF-1948256, CNS-195599, and CNS-1955997, the Air Force Office of Scientific Research under grant FA9550-23-1-0068, and the Army Research Laboratory under grants W911NF-19-10217 and W911NF-17-1-0094.